### v1

# Tint Your Models Task-wise for Improved Multi-task Model Merging

Abstract
Traditional model merging methods for multi-task learning (MTL) address task conflicts with straightforward strategies such as weight averaging, sign consensus, or minimal test-time adjustments. This presumably counts on the assumption that a merged encoder still retains abundant task knowledge from individual encoders, implying that its shared representation is sufficiently general across tasks. However, our insight is that adding just a single trainable task-specific layer further can bring striking performance gains, as demonstrated by our pilot study. Motivated by this finding, we propose Model Tinting , a new test-time approach that introduces a single task-specific layer for each task as trainable adjustments. Our method jointly trains merging coefficients and task-specific layers, which effectively reduces task conflicts with minimal additional costs. Additionally, we propose a sampling method that utilizes the difference in confidence levels of both merged and individual encoders. Extensive experiments demonstrate our method’s effectiveness, which achieves state-of-the-art performance across both computer vision and natural language processing tasks and significantly surpasses prior works. Our code is available at https://github.com/AIM-SKKU/ModelTinting.
1 Introduction
Multi-task learning (MTL) [67] aims to develop models capable of handling multiple tasks simultaneously, enhancing efficiency and enabling knowledge transfer across tasks. Recent advances have facilitated the merging of independently fine-tuned models into a unified framework, significantly improving performance. Inspired by weight-space averaging [59, 58], Task Arithmetic [19], which formalizes task-specific knowledge as task vectors — the difference between pre-trained and fine-tuned models — opened new horizons in MTL.
However, we argue that negative transfer between tasks, which is a challenge in traditional MTL, still persists in model merging. To mitigate task conflicts arising from model merging, recent methods [63, 10, 66, 18] have employed predefined strategies focused on reducing parameter interference in shared representations. While effective, these methods involve ad-hoc procedures and focus entirely on shared representations to capture task-specific characteristics. Relying solely on such fixed strategies could limit the preservation of task-specific knowledge, creating a bottleneck in model merging.
This raises two fundamental questions: 1) Is relying solely on predefined strategies using shared representations sufficient to preserve task-specific knowledge? 2) Is there an effective way to capture task-specific features without relying on complicated training methods? To answer these questions, we start with a pilot study revealing that pairing a single respective task-specific layer with different task encoders trained on different tasks consistently enhances cross-task performance. This suggests that adding a minimal trainable layer (e.g., the classifier head) that captures task-specific features complements the task-agnostic knowledge in shared representations. This finding became the foundation of our proposed method.
In this paper, we propose a novel test-time MTL merging method dubbed Model Tinting that introduces respective task-specific layers through minimal learnable adjustments. By refining each layer to focus on its respective knowledge, our approach captures distinct task details, allowing the merged model to extract relevant features while retaining shared representations. Unlike previous methods [64, 50, 65], the proposed method focuses on enhancing task-specific knowledge at the output level. Our approach jointly trains merging coefficients and task-specific layers without requiring explicit costly optimizations, as a single coefficient set and layer are assigned to each task. To prevent the merged model from adopting incorrect predictions from individual models, we further propose Model Tinting++ . This method introduces a sampling strategy that excludes instances where individual models are incorrect, but the merged model predicts correctly based on confidence.
We demonstrate Model Tinting through extensive experiments, which achieve state-of-the-art performance across both vision and NLP domains (partially visualized in Figure 1) for different model scales and task configurations. We also provide some justifications for our design and some insights drawn from the empirical studies. Beyond the performance gains, our work demonstrates that effective multi-task learning does not require heavy fine-tuning that disrupts shared representations. Instead, task-specific knowledge can be naturally preserved and enhanced through minimal task-specific adaptations.
2 Related Work
2.1 Multi-task Learning
Multi-task Learning (MTL) enables models to perform multiple tasks simultaneously by leveraging shared knowledge [52, 3, 69]. However, MTL faces challenges like task interference and negative transfer. To address these issues, modularization techniques introduce task-specific modules or pathways, preserving unique task information and reducing interference [33, 36, 45]. Gradient-based methods balance tasks through normalization and align gradient directions to minimize conflicts [5, 67]. Task weighting approaches dynamically assign weights to balance the learning process and prevent any single task from dominating [24, 46, 30, 17, 16]. Knowledge distillation further enhances MTL by transferring insights from specialized models to a unified framework [12, 21, 62]. Despite these advancements, traditional MTL methods often require extensive labeled data and significant computational resources. This highlights the need for more efficient and scalable approaches such as model merging, which can consolidate independently fine-tuned models into a unified framework.
2.2 Model Merging for Multi-task Learning
A foundational work in model merging for MTL is Task Arithmetic [19], which defines task vectors as the difference between fine-tuned and pre-trained models, enabling the combination of weights across multiple tasks through arithmetic operations. However, this method faces two main issues: parameter interference and maintaining task-specific information to prevent a merged model from becoming overly generalized. Recent approaches tackle interference through parameter drop, redundancy reduction, and test-time adjustments. For instance, Ties-Merging [63] addresses redundant parameters and sign conflicts; DARE [66] randomly removes redundancies and rescales key parameters; EMR-Merging [18] uses masks and rescaling to retain critical values; AdaMerging [65] learns the importance of task vectors at test time.
To preserve more task-specific details, WEMoE [50] and Twin-Merging [32] utilize trained routers to adjust shared and task-specific components, while Representation Surgery [64] refines the representation space with test-time training on unlabeled data. Overall, existing model merging methods assume shared features alone can capture all task-specific characteristics, but this often fails, losing critical task information. Our proposed method addresses this by focusing on task-specific layers rather than just shared representations. By refining these layers while keeping the encoder shared, Model Tinting captures essential task-specific features often overlooked in shared layers. This approach optimizes merging coefficients and task-specific layers, minimizing conflicts and often surpassing individual model performance.
Figure 1: Performance comparison of model merging methods. Radar charts highlight Model Tinting’s superiority across (a) 20 vision tasks and (b) 8 NLP tasks, achieving state-of-the-art results.
3 Preliminary
3.1 Background
Problem definition. Multi-task Learning (MTL) [67] aims to train a single model to achieve strong performance across multiple tasks. Given the input $x^{k}\in\mathbb{R}^{d_{in}}$ for the kth task, a $L$-layer encoder $f(x^{k};\phi)$ with the weights $\phi=\{\phi^{1},\ldots,\phi^{L}\}$ outputs the representation $\hat{z}^{k}\in\mathbb{R}^{d_{out}}$. For each of the $K$ tasks, $\hat{z}^{k}$ is fed into a task-specific layer $g(\cdot\ ;\theta_{k})$, where $\theta_{k}$ represents weights unique to task $k$, to perform classification (e.g., the classification head in CLIP [43]). Each task output should align with the task nature, and an improved MTL model should provide higher accuracies across tasks. We employ two types of weights to address MTL: $\phi_{\text{pre}}$ and $\phi_{k}$, which denote a pre-trained weight and task-specific fine-tuned weights for task $k$, respectively. These weights form the basis of novel weight merging methods [58, 22] aimed at excelling in MTL.
Task Arithmetic [19] introduced task vectors for each task $k$, defined as $\tau_{k}=\phi_{k}-\phi_{\text{pre}}$, to average weights for MTL. The merged weight is represented by $\phi_{\text{MTL}}=\phi_{\text{pre}}+\lambda\cdot\sum_{k=1}^{K}\tau_{k}$, where $\lambda$ is a scaling hyperparameter for the task vectors, determined using held-out validation sets rather than being learned. With the merged weight and individual classifiers, a single merged model $f(x;\phi_{\text{MTL}})$ is capable of managing all tasks with decent performance.
AdaMerging [65] improved the Task Arithmetic by building a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task-wise and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$. Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing the objective: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for a test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient. Although AdaMerging effectively merges multi-task weights, we believe there is still room for improvement.
3.2 Pilot Study
The success of prior MTL merging methods [32, 66, 63, 65, 18] lies in reducing conflicts in the encoder’s representations across tasks. However, we argue that adopting the shared representations without task-specific knowledge may harm MTL performance due to the task-agnostic nature of the merged encoder [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the merged encoder becomes more likely to be task-agnostic..
Study outline. Our goal is to validate the claim and take it a step further by showing that the following solution works: a trained task-specific layer in a merged model could enhance task-specific capabilities in task-agnostic representation. Here, we use the classifier as the layer for analysis, as it is the most straightforward option. If our claim is correct, the merging methods [65, 64] could benefit performance-wise from minimal additional training.
Cross-task evaluation. Given that each $k$-th task has its own fine-tuned model, $g_{k}^{bs}\left(f(x^{k};\phi_{k});\theta_{k}^{bs}\right)$, which serves as our control group, where $g_{k}^{bs}(\cdot;\theta_{k}^{bs})$ denotes the base (or primitive) classifier in the pre-trained model (e.g., CLIP [43]). We define the cross-task performance, computed by using the visual encoder from another task $m$ in the form $g_{k}^{bs}\left(f(x^{k};\phi_{m});\theta_{k}^{bs}\right)$ for a given task $k$ (i.e., swapping the trained layer only with the original one). We emulate the optimization-based merging [65] to observe whether the trained weight would steer towards a task-specific direction. To this end, a weight-averaged encoder [19], $f(x^{k};\phi_{\text{MTL}})$ is first employed as the backbone to generate task-agnostic representations. Training is performed to minimize the following objective with respect to the designated layers ($g_{k}^{tr}$)’s weight ($\theta_{k}^{tr}$):
$\underset{\{\theta_{k}^{tr}\}}{\min}\sum_{k,(x^{k},y^{k})\in\mathcal{X}}\mathcal{L}_{k}(g_{k}^{tr}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{tr}),y^{k}),$ (1)
where $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$. This setup uses labeled samples to confirm the encoder’s maximal capability by utilizing ground-truth labels. We then compute cross-task scores by pairing each trained classifier $g_{k}^{tr}(\cdot;\theta_{k}^{tr})$ with individual encoders, which is formulated as $g_{k}^{tr}(f_{m}(x^{k};\phi_{m});\theta_{k}^{tr})$ for all $m\neq k$.
(a) Primitive classifiers
(b) Trained classifiers
Figure 2: Cross-task evaluations across diverse encoders and heads from different tasks. Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier. The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal. (a) Primitive (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders. (b) Our trained classifiers show mostly reduced performance loss (i.e., higher relative numbers). This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise.
Our interpretation. Figure 2 shows cross-task evaluations via individual visual encoders with their respective primitive classifiers; Figure 2 shows results by combining the individual visual encoders with the trained classifiers. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. For instance, integrating the MNIST classifier with the Cars encoder achieves 58.1% accuracy (Figure 2), surpassing the base classifier’s 51.8% (Figure 2). These results suggest that training the classifier on merged features for the MNIST task also facilitates the alignment of individual Cars features with the MNIST task. This demonstrates the trained layer’s enhanced capability to refine task-specific outputs from task-agnostic representations. We further conjecture that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels for queries in practice. Therefore, our method is designed with this constraint in mind so that it can work in real-world scenarios.
4 Method
This section introduces our method, inspired by the pilot study, which incorporates task-specific adaptation during weight merging to improve MTL performance significantly.
4.1 Model Tinting
Motivation. Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific capabilities, potentially boosting overall MTL performance. Motivated by this finding, we explore an effective merging method that selectively adapts a few layers in the merged encoder to each task during the merging process. We conjecture that this would merit improved MTL performance by injecting task-specific knowledge into the task-agnostic encoder, barely incurring extra training costs.
Leveraging task-specific knowledge at test-time. We design our method based on test-time approaches [65, 64], where ground-truth labels are unavailable. A straightforward approach would be to adopt the entropy [65] as the metric replacing any supervisory signals. Since the entropy roughly follows the cross-entropy’s trend [65, 38], we believe that the self-entropy may not fully align when data cardinality is insufficient. We empirically found that the entropy does not closely approximate the cross-entropy (with ground-truths) during training (see Table 1), so we identify an effective way of employing the cross-entropy.
Drawing from what we found in the pilot study, we mimic each task encoder’s knowledge to equip the classifier with task-specific knowledge. We define an objective to align the distribution of each fine-tuned individual model, $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}\left(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}}\right)|{x^{k}{\in}\mathcal{X}}\}$, with the merged model’s, $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}}\right)|{x^{k}{\in}\mathcal{X}}\}$ on test data $\mathcal{X}$. While some semi-supervised techniques [68, 51, 41, 1] could be employed, we opt for a simpler approach [29]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach. We provide the related study in Figure 6. using the cross-entropy loss. This optimization objective is defined as follows:
$\underset{\{\lambda_{k}^{l}\},\{\theta_{k}^{\text{tr}}\}}{\min}\;\sum_{k=1}^{K}\mathcal{L}_{k}\left(C_{k}^{\text{merged}},C_{k}^{\text{ft}}\right),$ (2)
where $\mathcal{L}_{k}$ denotes the loss for each task, with $g_{k}^{\text{bs}}$ representing the primitive task-specific layers, and $g_{k}^{\text{tr}}$ representing the learnable task-specific layers. Notably, in contrast to AdaMerging [65], our approach jointly optimizes both the merging coefficients and the task-specific layers rather than focusing solely on the merging coefficients.
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.66 0.72 0.80 0.86 0.92 0.92 1.00 0.56
After 0.32 -0.47 0.79 0.94 0.83 0.41 0.96 0.47
$\Delta$0.34 -1.19 -0.01 +0.08 -0.09 -0.51 -0.04 -0.09
Ours Before 0.80 0.86 0.97 1.00 0.98 0.99 1.00 0.84
After 0.82 0.88 0.99 1.00 0.99 1.00 1.00 0.88
$\Delta$ +0.02 +0.02 +0.02 +0.00 +0.01 +0.01 +0.00 +0.04
Table 1: Spearman correlation of losses with ground truth cross-entropy loss. We report the correlation coefficient for two merged weights: ‘before’ (initial merged weights) and ‘after’ (trained merged weights under the different losses). We evaluate two losses: ‘Entropy’ (the entropy minimization objective) and ‘Ours’ (our self-label-based cross-entropy loss).
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg.
Individual 0.6018 0.9404 0.8922 0.9063 0.9141 0.8720 0.9271 0.7906 0.8556
Weight Averaging 0.1396 0.6411 0.6936 0.3184 0.7536 0.4219 0.5870 0.5523 0.5134
RegMean [23] 0.3667 0.9060 0.7574 0.6268 0.8355 0.7002 0.8235 0.5848 0.7001
Task Arithmetic [19] 0.1878 0.8589 0.7990 0.7403 0.8378 0.5908 0.6967 0.6209 0.6665
Ties-Merging [63] 0.2048 0.8440 0.8113 0.5819 0.8570 0.6465 0.7481 0.4296 0.6404
EMR-Merging [18] 0.3996 0.9335 0.8627 0.8277 0.8972 0.8545 0.8957 0.7437 0.8018
[…]
Model Tinting
0.6018 0.9392 0.8922 0.9049 0.8606 0.8382 0.8955 0.7906 0.8404
Table 2: Multi-task performance across 8 NLP tasks. We present a performance comparison of merging the RoBERTa models fine-tuned on 8 tasks in GLUE. $\dagger$ denotes our reproduced results for a fair comparison.
Loss correlation. Here, we justify our design choice as represented in the loss in Eq. (2). Inspired by the analysis [65], we demonstrate the effectiveness of using cross-entropy with model-generated labels by computing the Spearman correlation between cross-entropy losses based on individual model signals and ground-truths; the correlation of AdaMerging’s entropy loss versus the ground-truth-based cross-entropy loss is also computed. In addition, we further examine the correlation before and after training to see how training affects the correlation.
Table 1 reports the correlation coefficients for each task. We use the weight-averaged model [19] to provide baseline values in the ‘before’ rows, while ‘after’ denotes the correlations measured using a trained merged model. We observe that the entropy loss initially shares the correlations with the supervised cross-entropy loss, and the correlation mostly dropped after training across diverse tasks (i.e., 1 out of 8), particularly on challenging ones like Cars, even showed a negative correlation. This suggests that entropy minimization may not fully follow the true supervisory signals for MTL [^3] As also observed in prior studies [65, 38], entropy sometimes diverges further from the supervised cross-entropy for certain sample distributions.; the resulting merged model does not appear to be trained under ground-truth supervisory signals.
In contrast, our model-generated cross-entropy loss initially shows higher correlations and maintains a striking, strong positive correlation after training. Intriguingly, the correlation strengthens after training, which suggests that the merged model becomes more aligned with supervisory signals, likely due to its improvement under appropriately emulated signals. Therefore, our approach effectively trains a merged model that can better follow the ground-truth signals at test time when labels are inaccessible.
4.2 Model Tinting++
Improving Model Tinting. Model-generated supervisory signals are crucial for our method, and enhancing these signals could yield further improvements. Our insight is to exclude undesirable samples during training where the merged model is correct while individual models are incorrect in that training with such signals may not always be beneficial. How many such samples exist, and what is their impact? Figure 5 shows the extent to which the number of such cases is not negligible. Another question could be: how can we identify these samples for exclusion?
First, we presume only confident samples for individual models would be meaningful in training. Additionally, we refer to the literature [15], which suggests that a calibrated model achieves robustness by mitigating overconfidence; we similarly assume that a merged model would follow the trend and less-confident samples would be helpful [38]. Since quantifying confidence is challenging and introducing a new confidence threshold is undesirable, we sample based on the confidence trends in both individual and merged models. Specifically, we use the top-1 probability as the metric and filter out the instances when the merged model’s top-1 probability for a given prediction is higher than that of the individual model.
This approach can be understood as selecting only the samples that could calibrate the merged model while individual models provide reliable supervisory signals for training. This ensures that the resultant merged model becomes more robust, as similarly discussed in [39]; thereby, we expect improved MTL performance across different tasks. Our experimental results will demonstrate the improvements achieved by the proposed method.
Dicussions on implementation. Previous methods [65, 50] sample a batch for each task, pass each batch through a shared backbone, and then process it through the corresponding task-specific head, which is typical. The losses from each task are then combined, and a single gradient update is performed after completing the forward pass for all tasks. However, this approach becomes less efficient as the number of tasks or model size increases.
To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one. Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach. While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6. Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%). We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [26] by training tasks sequentially rather than simultaneously.
Method SUN Cars RES. Euro SVHN GTS. MNIST DTD Avg.
Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 48.0
Individual 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
Traditional MTL 73.9 74.4 93.9 98.2 95.8 98.9 99.5 77.9 88.9
Weight Averaging 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1 65.8
Fisher Merging [35] 68.6 69.2 70.7 66.4 72.9 51.1 87.9 59.9 68.3
RegMean [23] 65.3 63.5 75.6 78.6 78.1 67.4 93.7 52.0 71.8
Task Arithmetic [19] 55.2 54.9 66.7 78.9 80.2 69.7 97.3 50.4 69.1
[…]
Ties Merging [63] 65.0 64.4 74.8 77.4 81.2 69.3 96.5 54.5 72.9
[…]
Concrete TA [49] 62.5 61.1 76.0 95.7 91.0 81.9 98.5 51.9 77.3
Concrete AM [49] 67.8 70.0 87.5 96.0 91.6 96.7 98.7 63.8 84.0
AdaMerging++ [65] 66.6 68.3 82.2 94.2 89.6 89.0 98.3 60.6 81.1
Ada. w/ Surgery [64] 71.2 72.0 92.3 99.0 92.2 97.9 99.0 76.1 87.5
EMR-Merging [18] 75.2 72.8 93.5 99.5 96.9 98.1 99.6 74.4 88.7
WEMoE [50] 74.1 77.4 93.7 99.1 96.2 98.9 99.6 76.4 89.4
Model Tinting
73.7 77.6 94.7 99.0 95.4 98.2 99.1 79.1 89.6
Model Tinting++
74.3 79.3 94.8 99.0 95.7 98.5 99.2 80.2 90.1
(a) On merging ViT-B/32 models
Method SUN Cars RES. Euro SVHN GTS. MNIST DTD Avg.
Pretrained 66.8 77.7 71.0 59.9 58.4 50.5 76.3 55.3 64.5
Individual 82.3 92.4 97.4 100 98.1 99.2 99.7 84.1 94.2
Traditional MTL 80.8 90.6 96.3 96.3 97.6 99.1 99.6 84.4 93.5
Weight Averaging 72.1 81.6 82.6 91.9 78.2 70.7 97.1 62.8 79.6
Fisher Merging [35] 69.2 88.6 87.5 93.5 80.6 74.8 93.3 70.0 82.2
RegMean [23] 73.3 81.8 86.1 97.0 88.0 84.2 98.5 60.8 83.7
Task Arithmetic [19] 73.9 82.1 86.6 94.1 87.9 86.7 98.9 65.6 84.5
[…]
Ties Merging [63] 76.5 85.0 89.3 95.7 90.3 83.3 99.0 68.8 86.0
[…]
Concrete TA [49] 74.6 86.2 89.0 96.7 93.6 93.4 99.1 66.9 87.4
Concrete AM [49] 77.8 91.2 92.1 97.0 94.4 97.9 99.0 79.5 91.1
AdaMerging++ [65] 79.4 90.3 91.6 97.4 93.4 97.5 99.0 79.2 91.0
Ada. w/ Surgery [64] 80.3 90.8 94.3 98.2 94.1 98.7 99.2 82.5 92.3
EMR-Merging [18] 83.2 90.7 96.8 99.7 97.9 99.1 99.7 82.7 93.7
WEMoE [50] 81.4 92.6 95.4 99.4 97.7 99.3 99.7 83.7 93.6
Model Tinting
81.4 92.6 96.9 99.4 97.6 99.0 99.4 84.2 93.8
Model Tinting++
81.2 92.3 97.0 99.7 97.3 99.1 99.4 84.8 93.9
(b) On merging ViT-L/14 models
Table 3: Multi-task performance comparison across 8 vision tasks. We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models.
5 Experiment
5.1 Experimental Setup
Datasets and metrics. We follow the standard setups in DARE [66] and EMR-Merging [18] for NLP experiments: employing fine-tuned weights for 8 tasks from the GLUE benchmark [54]. Evaluation on each task is performed with task-specific metrics: Matthews correlation for CoLA [56], Pearson and Spearman correlations for STS-B [4], and accuracy for SST-2 [47], MRPC [11], QQP [20], MNLI [57], QNLI [44], and RTE [13]. For vision tasks, we follow the standard Task Arithmetic [19] and AdaMerging [65] setups, using fine-tuned weights on 8 standard image classification tasks. Furthermore, we extend our vision task experiments with 12 more tasks reaching the overall experimental scope to 20 tasks, following [55]: CIFAR-10 [28], CIFAR-100 [28], FER2013 [14], Flowers102 (Flowers) [37], Oxford-IIIT Pet (Pet) [40], PCAM [53], STL10 [8], EMNIST [9], FashionMNIST (FMNIST) [60], Food101[2], KMNIST [7], and Rendered SST-2 (R-SST) [42]. All tasks use the accuracy metric. Additionally, we provide results for a selected subset of 14 out of the 20 tasks in Appendix.
Models. We employ RoBERTa-base [31] for the NLP tasks in our experiments. We use the publicly released weights [^4] https://github.com/harveyhuang18/EMR_Merging by the prior work [18]. For a fair comparison, we run experiments for TWIN-Merging [32] and our method using fine-tuned weights from EMR-Merging. Our method applies Mean Squared Error (MSE) loss to STS-B instead.
For vision experiments, we utilize the pre-trained CLIP [43] weights [^5] https://github.com/mlfoundations/open_clip, specifically ViT-{B/32, B/16, L/14}. Most analyses and ablation studies are performed with ViT-B/32 for comparison. For a fair comparison across the 8 tasks, we use the fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors publicly released by Task Arithmetic [19]. In experiments with 14 and 20 vision tasks, CLIP ViT-B/32 is fine-tuned based on the configuration used in Task Arithmetic to obtain individual models. Due to the lack of 14 and 20 tasks experiments in literature, we reproduce Task Arithmetic, MagMax [34], Ties Merging [63], Consensus TA [55], AdaMerging [65], Surgery [64], and EMR-Merging [18]. More detailed experimental setups are provided in Appendix.
5.2 Main Results
Merging 8 NLP tasks. As shown in Table 2, our Model Tinting largely improved over the best method on average. While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks. We argue that this reveals the limitations of methods that only handle shared representations and shows the superiority of our method that effectively manages task-specific information. For further details, please refer to Appendix.
Merging 8 vision tasks. Table 3 and Table 3 demonstrate that our approach surpasses traditional MTL methods and achieves average performance improvements of 0.7%p and 0.2%p on ViT-B/32 and ViT-L/14, respectively, compared to the previous state-of-the-art merging methods. (For ViT-B/16, we observe an improvement of 0.7%p; more details are provided in Appendix) Additionally, our method achieves performance close to the individual models, even surpassing them on some tasks. When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal. This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
Merging 20 vision tasks. In Table 4, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance. Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic. In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p. Notably, for several tasks, our method even exceeds the performance of individual models. With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD CIFAR100 FER2013
[…]
Task $\dagger$ [19] 61.8 53.0 61.9 57.5 49.8 44.6 77.9 45.6 65.4 41.4
[…]
Model Tinting
72.9 76.1 94.5 98.7 92.4 97.4 99.1 79.5 84.7 69.2
Model Tinting++
73.5 78.6 94.5 99.2 93.3 97.2 99.3 79.7 83.9 68.5
Method Flowers Pet PCAM STL10 CIFAR10 EMNIST FMNIST Food101 KMNIST R-SST2
[…]
Task $\dagger$ 63.1 86.0 65.8 94.4 91.5 39.6 73.9 72.1 12.2 57.8
[…]
Model Tinting
90.9 91.9 82.1 97.8 96.0 98.3 90.6 85.5 95.3 75.2
Model Tinting++
92.0 92.4 84.7 98.0 96.0 98.8 91.2 85.2 96.3 76.4
Method Avg.
[…]
Task $\dagger$ 60.8
[…]
Model Tinting
88.4
Model Tinting++
88.9
Table 4: Task-specific and average performance across 20 vision tasks. We use ViT-B/32 for experiments. $\dagger$ denotes reproduced results.
5.3 Emprical Analyses
Sparsity analysis of merging coefficients. Inspired by the literature [66, 63, 10] in which redundant parameters degrade performance due to task-specific parameters’ conflicts, we analyze the learned merging coefficients to observe this effect in our method. We explore sparsity in the merging coefficients after training them alone and jointly with the classifier. Figure 3 shows both configurations sequentially and demonstrates that the joint training enjoys striking sparsity. We find 37.2% versus 55.9% of the coefficients are concentrated near zero (i.e., smaller than 1e-5) with the deviated average accuracy of 83.4% versus 89.6% for each configuration. This suggests that training a task-specific layer, like the classifier, along with merging coefficients, reduces task interference by pruning unnecessary parameters, leading to improved accuracy.
(a) Training coefficients only
(b) Joint training
Figure 3: Impact of joint training on coefficient sparsity. We visualize the merging coefficients to display the sparsity in weights. Joint training of merging coefficients with a task-specific layer promotes sparsity strikingly, with 55.9% of coefficients near zero compared to 37.2% when training coefficients only.
Can inner layers serve as task-specific adjustments? One might wonder if our observed improvements are unique to training just the classifier. We disprove this speculation by evaluating other layers beyond the classifier. Specifically, we train each of the 12 transformer layers in a merged encoder, which is performed under a similar setup in the pilot study. Figure 4 shows transformer layers generally achieve performance matched to training the classifier (see the details in Appendix). This suggests that training a single internal transformer layer can do similarly to what the classifier does. Furthermore, identifying the optimal layer could lead to additional performance improvements. Meanwhile, when we train either all layers or groups of layers — early layers (from layers 0 to 5-th) or late layers (from layers 6 to 11-th) together — performance drops significantly; we believe that training many layers at once in MTL scenarios diminishes the task-agnostic knowledge inherently stored in a pre-trained model.
Figure 4: Impact of training different task-specific layers. Performance comparison of merged models when training: 1) a single inner layer, 2) late layers, 3) early layers, and 4) only the classifier. Training a single layer matches the performance of training only the classifier, while training multiple layers reduces accuracy.
(a) Prediction discrepancy
(b) Overall difference
Figure 5: Prediction discrepancies between the merged model and individual model. (a) The upper bars indicate predictions correctly classified by individual models but misclassified by the merged model; the lower (hatched) bars indicate the opposite. A lower upper portion and a higher lower portion indicate a positive impact on the merged model performance. (b) represents the overall difference between the upper and lower bars.
Assesing merging methods via prediction discrepancy. Here, we assess diverse merged models with their sample prediction correctness with respect to the individual models that often serve as the performance upper bound in MTL due to higher accuracy. Figure 5 displays two sets of bar graphs. First, the upper bars represent samples correctly predicted by the individual models but incorrectly by the merged model, where a lower one aligns more with the individual models. Our method achieves the smallest discrepancy, significantly outperforming AdaMerging and Surgery. The lower bars in Figure 5 represent samples misclassified by the individual models but correctly classified by the merged model, representing unexpected gains of the merged model. A higher one here benefits from the merged model’s predictions beyond those of the individual model.
Figure 5 illustrates the overall impact by computing the difference between the penalties and gains shown in the upper and lower bars of Figure 5. First, our improved method presented in Section 4.2 successfully enhances this metric, where a lower one is better. We believe our methods effectively balance preserving the strengths of the individual models while capturing unexpected gains, leading to superior overall performance. We observe that while closely related methods like AdaMerging [65] (using entropy minimization without constraints) and Surgery [64] (performs post-AdaMerging to preserve individual model predictions) also enjoy some unexpected gains, the penalties observed in the upper bars outweigh the benefits, which incur net decreases in the overall metric.
(a) On supervisory model
(b) Training robustness
Figure 6: More analyses with merging coefficients. (a) Refined supervisory models merged under various coefficients are tested. Our design choice – using the unmerged individual models – performs near-optimally. (b) MTL accuracy confirms the initialization robustness of our method across different initial merging coefficients; related results are illustrated in Figure 7.
More studies with merging coefficients. We study employing individual model predictions without any refinement is a sensible choice for supervisory signals. We refine predictions by merging individual models with a pre-trained model, which is expected to yield more precise predictions. Specifically, we handle each individual model as a pre-trained model with an added task vector scaled by a merging coefficient, ranging from 0 (pre-trained model) to 1 (individual model). We experiment across these settings, and the results are shown in Figure 6, where performance consistently improved as predictions approached the individual model. This indicates that guiding the classifier with a task-agnostic pre-trained model limits its ability to capture task-specific details effectively. Though a combined model (at the coefficient 0.8) performs the best, which is about 89.9%, our design choice of using individual models alone achieves a near-best result. Finding the optimal coefficient would require cumbersome hyperparameter tuning, so using the existing individual models directly is a practical option. Efficiently searching for the coefficient would be a potential research direction in the future.
Initialization robustness. Our method benefits from the robustness of the merging coefficients’ initialization. To highlight this, we present training curves under different initialization settings and compare them with those of other model merging methods. As shown in Figure 6, when only the classifier is trained, the performance is sensitive to initialization; however, when both the classifier and the coefficients are trained together, the method demonstrates substantial robustness. Figure 7 shows that when training a Surgery model or the classifier only, starting from the AdaMerging initialization yields higher performance than starting from the Task Arithmetic initialization, which suggests the sensitivity to initialization.
Figure 7: Impact of initializations. We test two different approaches: fixed-value initialization from Task Arithmetic (dashed lines) and 2) learned-value one in AdaMerging (solid lines). The learning trend diverges with what to learn, but training both coefficients and task-specific layers is not affected by the initializations.
6 Conclusion
In this paper, we have introduced Model Tinting , a novel approach to multi-task model merging that enhances task-specific knowledge through adaptively learning task-specific layers, departing from traditional methods that rely on shared representations. Model Tinting++ further involved a novel confidence-based sampling method, which has proven favorable. Extensive experiments demonstrated state-of-the-art performance across various domains, including vision and NLP, with particularly significant improvements on the challenging 20-task vision benchmark. Our empirical studies further justified that our design choice of incorporating just a single task-wise learned layer is sensible. Finally, we discovered additional insights: 1) inner layers all play similar roles in task-specific adaptations; 2) training task-specific layers naturally promotes sparsity in merging coefficients; 3) our joint training is insensitive to initializations with improved results. We believe our study could be a foundation for future research in model merging.
Appendix
This Appendix provides an overview of our experimental setup and baselines, further experimental results across tasks, and additional empirical analyses. The detailed descriptions of each section are summarized as follows:
  • Appendix A: Outline of Model Tinting with introducing an illustrative figure;
  • Appendix B: Detailed experimental setup and hyperparameter configurations;
  • Appendix C: Additional results for NLP and vision tasks;
  • Appendix D: Further analyses on task-specific layers, loss functions, convergence, and sparsity.
Appendix A Overview of Model Tinting
We give a brief overview of our proposed method Model Tinting , a test-time approach for merging multiple fine-tuned models by introducing a single trainable layer for each task. Figure A provides an illustrative overview of our method. At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using merging coefficients. During this process, a trainable task-specific layer ($\phi_{\text{tint}}^{l}$) is introduced at a chosen transformer layer $l$, capturing task-specific information from the task-agnostic representations of the merged encoder. We argued that any layer can be picked for task adjustments before merging, and this is further supported by the graph in Figure A. The results demonstrate that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline. This highlights the robustness and effectiveness of task-specific adjustments in enhancing task-agnostic representations, with more details provided in Appendix D.1. We also proposed Model Tinting++ , which applies a strategy to filter noisy predictions from individual models.
Appendix B Experimental Setup
B.1 Training Details
Main experiments. We initialize the layer-wise merging coefficients for all layers to 0.3 by default, following the values used in the previous method [65]. However, for the 14 and 20 vision tasks, the coefficients are initially set to 0.1, as using 0.3 incurred significantly lower Task Arithmetic [19] performance for these tasks, presumably due to the increased number of tasks. We use the Adam [25] optimizer with momentum parameters (0.9, 0.999) to update the coefficients and the classifier. For vision tasks, the learning rate is set to 0.01 when training the classifier and 0.001 when training only the merging coefficients. For NLP tasks, a learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated. Consistent with prior works [65, 50, 64], vision tasks are trained with 16,000 test samples per task over 500 iterations, using a batch size of 32. Similarly, NLP tasks are trained with 32,000 samples per task over 1,000 iterations, also with a batch size of 32. For each task, the loss is computed, and the model parameters are updated immediately following the forward pass for each batch. To enable sequential processing, the order of tasks is randomized.
Pilot study. The aforementioned training details are applied almost identically to the experiments in the pilot study presented in our main paper. As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic-merged weights for the 8 vision tasks using ViT-B/32. A difference is that, instead of using the test dataset, the training dataset is used in a supervised manner to produce the results closer to the upper bound, in which the difference stands out more. We train the models for only one epoch.
B.2 Baseline Details
Figure A: Illustration of   Model Tinting . Our method improves multi-task learning by introducing a trainable, task-specific layer during model merging. Our approach (red) consistently achieves better results than the baseline without tinting (green) across all layers.
We compare our method against a diverse set of baselines, ranging from simple merging strategies to advanced methods leveraging task-specific or shared representations.
Pretrained indicates a model that predicts multiple tasks without additional fine-tuning for task-specific requirements. However, the absence of task-specific information for downstream tasks generally leads to poor performance.
Individual refers to the fine-tuning of individual pre-trained models for each task. Since there is no interference between tasks, it has been regarded as the upper bound of task-specific performance.
Traditional MTL trains a multi-task model by combining the original training data for all tasks. While effective, this approach depends on access to raw data and labels for each task, which is not always practical.
Weight Averaging merges multiple individual models by directly averaging their parameters to create a single model for multi-task learning. Although simple, this method lacks task-specific adjustments.
Fisher Merging [35] uses the Fisher information matrix to estimate the importance of each parameter. Model parameters are merged based on their relative contributions.
RegMean [23] merges models by aligning the inputs of linear layers to minimize the L2 distance between the individual models and the merged model.
Task Arithmetic [19] defines the difference between the fine-tuned and pre-trained model parameters as a task vector. By combining multiple task vectors and adding them to the pre-trained model, it enables multi-task learning.
MagMax [34] merges task vectors [19] by selecting the parameter with the largest magnitude for each position, consolidating knowledge into a single model without retaining task-specific data.
Ties Merging [63] highlight the importance of addressing interference in task arithmetic-based merging. It involves removing redundant parameters from the task vector and resolving parameter sign conflicts.
Consensus TA [55] enhances Task Arithmetic [19] by using Consensus Merging, which retains only weights beneficial to multiple tasks while eliminating irrelevant or task-specific weights. This process uses task-specific binary masks to identify relevant weights and forms a consensus mask to minimize task interference.
Concrete TA [49] utilizes a discrete binary mask to identify a low-dimensional subspace shared across tasks, where Task Arithmetic [19] is applied for model merging.
Concrete AM [49] extends Concrete TA by incorporating the training of the merging coefficients [65] during the merging process.
AdaMerging [65] adaptively learns merging coefficients at the task or layer level by minimizing the entropy of predictions on unlabeled test data.
Representation Surgery [64] reduces representation bias by training a task-specific module that aligns the merged model’s features with those of the individual models.
EMR-Merging [18] involves three steps: Elect, Mask, and Rescale-Merging. These steps select key parameters to form a unified model, apply task-specific masks for each task, and adjust scales to achieve better performance.
WEMoE [50] utilizes a Mixture of Experts (MoE) module to dynamically separate and integrate shared and task-specific knowledge based on input samples. By training the router on unlabeled test data, it optimizes routing weights and improves task-specific flexibility and performance.
Twin-Merging [32] employs the MoE module but introduces a preprocessing step where task-specific knowledge is compressed into sparse vectors via Singular Value Decomposition. During inference, the router integrates shared and task-specific components dynamically based on input.
Method SUN Cars RES. Euro SVH. GTS. MNI. DTD C100 FER Flower Pet PCAM STL Avg.
Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 64.2 39.0 66.3 87.4 60.6 97.1 57.1
[…]
Model Tinting
73.3 77.1 94.4 98.8 94.2 98.0 98.8 79.0 84.9 70.6 90.9 91.7 85.4 97.9 88.2
Model Tinting
73.7 78.5 94.2 99.3 94.6 97.8 98.9 80.2 84.4 70.5 92.1 92.2 84.8 98.0 88.5
Table A: Multi-task performance when merging ViT-B/32 models on 14 tasks. $\dagger$ denotes our reproduced results.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Pretrained 63.8 64.6 65.7 54.5 52.0 43.3 51.7 45.1 55.0
Individual 81.8 86.8 96.9 99.7 97.8 99.1 99.7 82.0 92.9
Weight Averaging 67.7 70.0 75.3 79.5 74.9 60.1 94.4 43.8 70.7
Fisher Merging [35] 68.5 69.9 75.2 80.4 73.2 61.2 94.5 50.7 71.7
RegMean [23] 69.1 71.6 77.6 88.8 83.7 70.2 96.9 54.6 76.6
Task Arithmetic [19] 61.1 65.9 74.0 76.2 88.0 73.9 98.4 53.0 73.8
[…]
Ties Merging [63] 69.1 72.5 80.5 84.0 85.0 71.5 98.1 54.9 77.0
[…]
AdaMerging++ [65] 71.8 80.8 84.1 94.3 91.9 94.5 98.7 69.8 85.7
Ada. w/ Surgery [64] 73.6 81.5 90.4 98.5 93.2 97.4 98.9 77.0 88.8
EMR-Merging [18] 78.6 82.6 95.5 99.2 97.6 98.8 99.6 78.3 91.3
Model Tinting
76.7 87.5 95.4 98.6 96.2 98.7 99.1 82.4 91.8
Model Tinting
77.4 87.0 95.6 98.6 97.0 98.9 99.2 82.7 92.0
Table B: Multi-task performance when merging ViT-B/16 models on 8 tasks.
Appendix C More Experimental Results
Figure B: Multi-task performance when merging RoBERTa models on 8 tasks. Yellow indicates training only coefficients, green indicates training only the classifier, and blue indicates training both in our method.
C.1 NLP Tasks
We conduct experiments to investigate how the training of the classifier and merging coefficients, respectively, affects the performance of NLP tasks in Model Tinting , as presented in Table 2 of the main paper. Figure B shows the results of merging RoBERTa [31] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss. As mentioned in the main paper, evaluation metrics differ across tasks: the Matthews correlation coefficient is used for CoLA [56], the average of Pearson and Spearman correlations is applied to STS-B [4], and the accuracy is used for all other tasks [47, 11, 20, 57, 44, 13]. The approach where both the classifier and coefficients are trained simultaneously corresponds to Model Tinting . To examine the role of classifier training in NLP tasks, ablation experiments are conducted by removing specific components. The results show that training only the coefficients leads to the lowest performance while training only the classifier achieves a relatively high performance. Furthermore, training both the classifier and coefficients together demonstrated a complementary effect, achieving the highest performance.
C.2 Vision Tasks
Table A presents the results of merging ViT-B/32 across 14 tasks. Our method outperforms the current state-of-the-art, EMR-Merging [18], by an average of 2.4%p. Notably, Model Tinting outperforms individual models on datasets such as Cars [27], DTD [6], Flowers [37], Pets [40], and STL10 [8], demonstrating its robustness and effectiveness across a wide range of tasks and datasets, regardless of the number of tasks. Table B shows the results of merging ViT-B/16 across 8 vision tasks. While other methods exhibit significant performance gaps compared to individual models, particularly on datasets like Cars and DTD, Model Tinting consistently achieves performance close to that of individual models.
Figure C: Impact of training different task-specific layers with merging coefficients. We compare average performance for merged models with different trainable task-specific layers while other layers update only their merging coefficients.
Appendix D More Empirical Analyses
D.1 On Task-specific Layers
Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
0 66.2 71.9 92.2 99.0 94.2 96.5 99.1 77.0 87.0
1 65.1 69.9 92.4 98.8 94.2 97.4 99.0 77.6 86.8
2 64.4 69.8 92.3 99.0 93.8 97.3 98.6 77.4 86.6
3 64.3 68.3 92.2 99.1 92.8 98.3 99.1 78.2 86.5
4 64.8 69.5 93.3 98.6 94.6 98.1 99.2 78.2 87.0
5 64.6 70.7 91.3 99.0 94.8 98.3 99.1 78.8 87.1
6 67.4 70.7 94.4 98.9 92.3 98.3 99.2 79.1 87.5
7 67.1 66.8 94.4 99.7 94.8 99.2 99.3 79.2 87.6
8 66.1 69.8 93.6 99.0 94.2 98.2 99.5 78.6 87.4
9 68.2 66.5 94.6 99.6 91.9 98.5 99.2 79.5 87.3
10 68.3 73.0 94.0 99.6 92.5 97.9 99.1 79.4 88.0
11 68.2 69.0 94.2 99.4 92.3 97.5 99.0 79.3 87.4
Classifier 69.9 73.8 92.2 98.6 89.5 97.5 99.0 78.8 87.4
Late 50.7 63.4 89.3 99.6 93.4 95.9 99.3 75.5 83.4
Early 42.5 54.2 85.0 93.8 93.2 96.3 98.2 76.0 79.9
All 25.5 31.2 82.7 97.0 92.4 96.2 97.5 72.7 74.4
(a) Training layers.
Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
0 73.5 76.8 95.2 99.7 95.9 98.4 99.4 79.4 89.8
1 73.5 76.8 95.6 99.9 96.2 98.6 99.4 79.4 89.9
2 73.9 76.8 95.4 99.8 96.2 98.6 99.3 79.5 89.9
3 73.8 76.7 95.4 99.9 96.5 99.0 99.4 79.4 90.0
4 74.1 77.0 95.5 99.8 96.5 98.6 99.4 79.4 90.0
5 73.9 76.7 95.6 99.9 96.3 98.3 99.4 79.4 89.9
6 74.1 77.1 95.5 99.7 96.4 98.8 99.5 79.4 90.1
7 74.1 75.6 95.3 99.8 95.9 98.6 99.6 79.4 89.8
8 74.1 77.3 95.6 99.9 96.1 98.3 99.2 79.4 90.0
9 74.5 77.3 95.7 99.7 96.0 98.4 99.4 79.4 90.0
10 74.5 77.2 95.5 99.8 96.3 98.5 99.5 79.4 90.1
11 74.3 76.7 94.7 99.8 95.8 98.2 99.3 79.4 89.8
Classifier 73.7 77.6 94.7 99.0 95.4 98.2 99.1 79.1 89.6
Late 69.6 73.1 94.2 99.7 96.0 98.6 98.4 79.1 88.6
Early 70.0 75.2 93.4 99.3 93.9 98.1 99.2 79.3 88.5
All 67.1 69.5 94.2 98.4 95.1 97.7 98.8 75.4 87.0
(b) Training layers with merging coefficients.
Table C: Performance details of merged ViT-B/32 models when training different layers for task-specific adjustments. We observe that consistently high performance can be achieved regardless of which task-specific single layer is trained. Note that the 0-th layer in the table refers to the first transformer layer.
Our proposed method, Model Tinting , introduces the concept of incorporating a task-specific layer to balance shared and task-specific representations during model merging. This approach allows effective adaptation while maintaining efficiency to address task conflicts in multi-task learning. Building on the results in Figure 4 of the main paper, which demonstrate the potential of training inner layers while keeping the merging coefficients fixed, we extend this analysis further.
Specifically, we explore whether performance remains robust when both merging coefficients and task-specific layers are trained jointly. ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers again. When a particular layer is marked as task-specific, it is split into unique versions for each task, resulting in eight task-specific layers. While these layers are trained exclusively on corresponding task data, the remaining layers update only the merging coefficients using data from all tasks.
Figure C shows that training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier. These findings highlight the minimal impact of task-specific layer choice on overall performance, reaffirming the robustness of our approach. Table C provides detailed task-level results, while Table C shows consistent average performance across layers, with specific layers excelling on certain datasets (e.g., the 11th layer on SUN397 [61] and the 7th layer on GTSRB [48]). Conversely, Table C demonstrates that merging coefficients adaptively balance representations across tasks, preventing biases even when other layers remain frozen.
D.2 On Loss Functions
Figure D: Comparison of training with various loss functions. The yellow bars represent the case where only the coefficients are trained, while the blue bars indicate the joint training of the classifier and coefficients. The dashed line corresponds to training with ground truth labels using cross-entropy loss, which is the upper bound.
(a) Before training (left: EM Merging   right: GT)
(b) After training (left: EM Merging   right: GT)
Figure E: t-SNE visualization results on the DTD dataset before (a) and after (b) entropy minimization training. The plots on the left use merged model predictions, while those on the right use ground truth labels. Training improves clustering, but prediction-ground truth mismatch remains.
In Figure D, we analyze the impact of various loss functions on aligning predictions between individual models and the merged model. When training only the coefficients, distance-based losses (e.g., L1, Smooth-L1, L2, and Cosine) demonstrate relatively lower performance compared to distribution-based losses (e.g., KL-divergence, JS-divergence, cross-entropy, and entropy). In contrast, when both the classifier and coefficients are trained jointly, most loss functions achieve strong performance under our method. However, entropy minimization, as explored in previous works [50, 65], leads to a significant decline in performance during joint training.
Notably, in Table 1 of the main paper, we observe that the entropy loss often fails to provide gradients in the correct direction due to its divergence from the ground truth (GT)-based loss. The table highlights that the cross-entropy loss, in contrast, aligns more closely with the GT, resulting in more accurate gradient directions. This highlights the importance of incorporating label information for guiding classifier training in the merged model. To further support this point, Figure E visualizes the t-SNE visualization results for the DTD dataset before and after entropy minimization training. While feature clustering improves after training, as shown in (b) compared to (a), a noticeable discrepancy remains between the predictions and the ground truth labels. These observations emphasize the importance of label guidance, even if incomplete, for achieving better alignment with the GT and improving the performance of the merged model.
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.73 0.82 0.84 0.86 0.96 0.91 1.00 0.61
After 0.38 0.31 0.94 0.92 0.88 0.96 0.98 0.38
$\Delta$0.35 -0.51 +0.1 +0.06 -0.08 +0.05 -0.02 -0.23
Ours Before 0.82 0.92 0.98 1.00 0.99 1.00 1.00 0.85
After 0.86 0.95 1.00 1.00 0.99 1.00 1.00 0.90
$\Delta$ +0.04 +0.03 +0.02 +0.00 +0.00 +0.00 +0.00 +0.05
(a) ViT-B/16
SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
Entropy Before 0.85 0.93 0.95 0.98 0.96 0.96 1.00 0.79
After 0.63 0.76 0.94 0.81 0.79 0.95 0.82 0.83
$\Delta$0.22 -0.17 -0.01 -0.17 -0.17 -0.01 -0.18 +0.04
Ours Before 0.87 0.96 0.98 1.00 0.98 1.00 1.00 0.89
After 0.89 0.97 1.00 1.00 0.99 1.00 1.00 0.91
$\Delta$ +0.02 +0.01 +0.02 +0.00 +0.01 +0.00 +0.00 +0.02
(b) ViT-L/14
Table D: Spearman correlation of losses with ground truth cross-entropy loss for (a) ViT-B/16 and (b) ViT-L/14. Values closer to 1 indicate that the corresponding loss function exhibits better alignment with the loss computed using the ground truth.
Loss correlation. We conduct further experiments to validate whether the loss correlation shown in Table 1 of the main paper remains consistent across different backbones (ViT-{B/16, L/14}). Table D shows the loss correlation results for 8 vision tasks using both the merged ViT-B/16 and ViT-L/14 models. When both the merging coefficients and the classifier are trained simultaneously, we observe a significant drop in correlation after training with entropy minimization. In contrast, the cross-entropy loss employed in our self-labeling approach maintains consistently high correlation across tasks. These findings are consistent with the results reported for ViT-B/32 in the main paper, suggesting that our self-labeling approach with the cross-entropy loss could be more effective than the entropy loss when training the classifier alongside the coefficients.
D.3 Convergence Analysis
Figure F illustrates the learning curve of Model Tinting applied to the 8 vision tasks over training iterations. In Figure F, we compare the performance of ViT-B/32, ViT-B/16, and ViT-L/14 models. The results demonstrate that our method achieves near-optimal performance within the first 100 iterations and converges quickly, regardless of model size. Notably, larger models like ViT-L/14 consistently achieve higher final accuracies, followed by ViT-B/16 and ViT-B/32, which highlights the advantages of model capacity in capturing task-specific knowledge more effectively. Figure F explores the impact of learning rates on the convergence of ViT-B/32. Except for the case with the lowest learning rate, the figure shows that our method converges quickly to a similarly high level of performance across most learning rates. These observations demonstrate the robustness and efficiency of Model Tinting across different model scales and hyperparameter settings.
(a) Comparison of average accuracy for ViT-B/32, ViT-B/16, and ViT-L/14 across training iterations.
(b) Average accuracy of ViT-B/32 across training iterations under different learning rates.
Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates.
(a) Training coefficients only
(b) Joint training
Figure G: Impact of joint training on coefficient sparsity on ViT-B/32.
(a) Training coefficients only
(b) Joint training
Figure H: Impact of joint training on coefficient sparsity on ViT-B/16.
(a) Training coefficients only
(b) Joint training
Figure I: Impact of joint training on coefficient sparsity on ViT-L/14.
D.4 Sparsity Visualization
The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging as shown in Figure 3. To confirm this trend across various backbones, Figure GH and I provide the layer-wise merging coefficients for ViT-B/32, ViT-B/16 and ViT-L/14. All models are optimized using the cross-entropy loss with the predictions of individual models as guidance. In these figures, (a) represents training only the coefficients, while (b) includes joint training of the classifier and coefficients. Joint training consistently increases the proportion of near-zero coefficients (from 35.3% to 56.6% for ViT-B/16 and from 23.3% to 55.0% for ViT-L/14). Correspondingly, the average accuracy also improves (from 86.8% to 91.8% for ViT-B/16 and from 91.5% to 93.8% for ViT-L/14). These results indicate that training the classifier complements sparsity in merging coefficients, effectively reducing task conflicts and enhancing performance.

### v2

@@
-Tint Your Models Task-wise for Improved Multi-task Model Merging
+Why Train Everything? Tint a Single Layer for Multi-task Model Merging
 Abstract
-Traditional model merging methods for multi-task learning (MTL) address task conflicts with straightforward strategies such as weight averaging, sign consensus, or minimal test-time adjustments.
-This presumably counts on the assumption that a merged encoder still retains abundant task knowledge from individual encoders, implying that its shared representation is sufficiently general across tasks.
-However, our insight is that adding just a single trainable task-specific layer further can bring striking performance gains, as demonstrated by our pilot study.
-Motivated by this finding, we propose Model Tinting , a new test-time approach that introduces a single task-specific layer for each task as trainable adjustments.
-Our method jointly trains merging coefficients and task-specific layers, which effectively reduces task conflicts with minimal additional costs.
-Additionally, we propose a sampling method that utilizes the difference in confidence levels of both merged and individual encoders.
-Extensive experiments demonstrate our method’s effectiveness, which achieves state-of-the-art performance across both computer vision and natural language processing tasks and significantly surpasses prior works.
+Model merging integrates independently fine-tuned models into a single multi-task model, offering a flexible alternative to joint training.
+However, many existing model merging methods introduce additional task-specific components, increasing complexity and requiring extra modifications.
+We propose Model Tinting , a lightweight yet highly effective approach that improves model merging by updating just a single layer, accounting for as low as 0.5% of total parameters.
+Our key observation is that explicit task-specific modules are not necessary; instead, subtle adjustments to a single layer can effectively capture task-specific variations within the merged model while maintaining generalization.
+We introduce a confidence-based filtering mechanism to alleviate the impact of unreliable predictions from individual models on the merged model.
+Extensive experiments across vision and NLP tasks demonstrate that Model Tinting achieves state-of-the-art performance, even in challenging dense prediction tasks.
 Our code is available at https://github.com/AIM-SKKU/ModelTinting.
 1 Introduction
-Multi-task learning (MTL) [67] aims to develop models capable of handling multiple tasks simultaneously, enhancing efficiency and enabling knowledge transfer across tasks.
-Recent advances have facilitated the merging of independently fine-tuned models into a unified framework, significantly improving performance.
-Inspired by weight-space averaging [59, 58], Task Arithmetic [19], which formalizes task-specific knowledge as task vectors — the difference between pre-trained and fine-tuned models — opened new horizons in MTL.
-However, we argue that negative transfer between tasks, which is a challenge in traditional MTL, still persists in model merging.
-To mitigate task conflicts arising from model merging, recent methods [63, 10, 66, 18] have employed predefined strategies focused on reducing parameter interference in shared representations.
-While effective, these methods involve ad-hoc procedures and focus entirely on shared representations to capture task-specific characteristics.
-Relying solely on such fixed strategies could limit the preservation of task-specific knowledge, creating a bottleneck in model merging.
-This raises two fundamental questions: 1) Is relying solely on predefined strategies using shared representations sufficient to preserve task-specific knowledge?
-2) Is there an effective way to capture task-specific features without relying on complicated training methods?
-To answer these questions, we start with a pilot study revealing that pairing a single respective task-specific layer with different task encoders trained on different tasks consistently enhances cross-task performance.
-This suggests that adding a minimal trainable layer (e.g., the classifier head) that captures task-specific features complements the task-agnostic knowledge in shared representations.
-This finding became the foundation of our proposed method.
-In this paper, we propose a novel test-time MTL merging method dubbed Model Tinting that introduces respective task-specific layers through minimal learnable adjustments.
-By refining each layer to focus on its respective knowledge, our approach captures distinct task details, allowing the merged model to extract relevant features while retaining shared representations.
-Unlike previous methods [64, 50, 65], the proposed method focuses on enhancing task-specific knowledge at the output level.
-Our approach jointly trains merging coefficients and task-specific layers without requiring explicit costly optimizations, as a single coefficient set and layer are assigned to each task.
-To prevent the merged model from adopting incorrect predictions from individual models, we further propose Model Tinting++ .
-This method introduces a sampling strategy that excludes instances where individual models are incorrect, but the merged model predicts correctly based on confidence.
-We demonstrate Model Tinting through extensive experiments, which achieve state-of-the-art performance across both vision and NLP domains (partially visualized in Figure 1) for different model scales and task configurations.
-We also provide some justifications for our design and some insights drawn from the empirical studies.
-Beyond the performance gains, our work demonstrates that effective multi-task learning does not require heavy fine-tuning that disrupts shared representations.
-Instead, task-specific knowledge can be naturally preserved and enhanced through minimal task-specific adaptations.
+Multi-task learning (MTL) [4] enables models to handle multiple tasks efficiently by facilitating knowledge transfer.
+Model merging [67, 22, 68] emerged as an alternative to conventional MTL [76, 40, 28, 20] by combining independently fine-tuned models into a single model.
+This approach allows for flexible adaptation, reduces computational overhead, and better preserves task-specific knowledge.
+A key challenge in model merging is balancing generalization and task specialization.
+On one hand, a merged model should retain shared knowledge across tasks [72, 74], but on the other hand, it must preserve task-specific adaptations that are crucial for individual task performance [73, 21].
+Figure 1: Performance comparison of model merging methods. Radar charts highlight Model Tinting’s superiority across (a) 20 vision tasks and (b) 8 NLP tasks, achieving state-of-the-art results.
+Early methods [23, 72, 75, 74, 71] primarily focused on constructing task-agnostic models by averaging weights or aligning parameters across models, assuming that similar models in weight space would naturally form a meaningful shared representation.
+However, this generalization-first strategy often led to a loss of task-specific information, reducing performance on individual tasks.
+This issue became more pronounced when merging models trained on diverse or conflicting tasks, as simple averaging often failed to preserve specialized adaptations essential for high accuracy.
+Recent approaches [21, 73, 58, 36] explored task-specific parameterization, with some incorporating additional components, in some cases, are trainable to enhance adaptation.
+These methods typically added task-specific modules or learnable merging coefficients, selectively adjusting how different tasks influence the final model.
+While this enhanced task adaptation, it also introduced challenges such as increased model complexity, higher computational costs in both training and inference, and scalability issues when applied to many tasks.
+Furthermore, many merging strategies still relied on a fixed shared model [23, 72, 75, 11, 36, 21, 73, 58], often formed by merging encoders without adaptation.
+As a result, the merged model struggled to retain fine-grained task distinctions and failed to effectively leverage task-relevant cues, leading to suboptimal performance in diverse task distributions (see Figure 1).
+These challenges raise two fundamental questions: 1) Can task-specific features be captured without introducing additional modules?
+2) Is a predefined task-agnostic backbone sufficient to retain task-specific knowledge effectively?
+In this paper, we propose Model Tinting , a simple yet effective model merging method that preserves task-specific knowledge with minimal overhead.
+Instead of introducing additional modules, our method maintains task-specific information through minimal trainable adjustments, updating a single layer that accounts for fewer than 0.5% of parameters in ViT-B/32.
+By jointly optimizing task-specific layers and merging coefficients, it refines both components under the guidance of individually fine-tuned models.
+To further enhance reliability, we introduce a confidence-based filtering mechanism that prevents the model from inheriting incorrect predictions from individual experts.
+By selectively removing unreliable supervision signals, this mechanism improves robustness and enables the merged model to sometimes surpass individual task models.
+Pub. Method Task-specific Adaptation No Extra Module Adaptive Shared Model
+ICLR 2023 Task Arithmetic [23]
+ICLR 2024 AdaMerging [74]
+ICML 2024 Surgery [73]
+ICML 2024 WEMoE [58]
+NeurIPS 2024 EMR-Merging [21]
+NeurIPS 2024 TWIN-Merging [36]
+Model Tinting
+Table 1: Comparison of recent model merging methods.
+Table 1 shows the main differences between existing methods and our method.
+Model Tinting requires no additional modules beyond a single trainable layer per task, making it lightweight and easy to deploy.
+Furthermore, we also adjust the task-agnostic model (i.e., shared model) to better leverage the effect of the task-specific layer, further enhancing adaptation for each task.
+Through extensive experiments across image classification, dense prediction, and NLP tasks, we demonstrate that Model Tinting achieves state-of-the-art performance across different model scales and task configurations.
+Notably, our method excels in dense prediction tasks, a challenging yet underexplored domain in prior model merging studies.
+Our approach achieves remarkably strong results, demonstrating its ability to handle complex task structures effectively.
+Our main contributions are as follows:
+  • We propose a lightweight MTL model merging method that preserves task-specific knowledge with minimal trainable adjustments, requiring no additional modules.
+  • To improve robustness, we introduce a confidence-based filtering mechanism that selectively removes unreliable supervision signals from individual models, preventing error propagation in the merged model.
+  • Our method achieves state-of-the-art performance across diverse tasks, including image classification, dense prediction, and NLP, particularly excelling in dense prediction, an underexplored area in model merging.
 2 Related Work
@@
 2.2 Model Merging for Multi-task Learning
-A foundational work in model merging for MTL is Task Arithmetic [19], which defines task vectors as the difference between fine-tuned and pre-trained models, enabling the combination of weights across multiple tasks through arithmetic operations.
-However, this method faces two main issues: parameter interference and maintaining task-specific information to prevent a merged model from becoming overly generalized.
-Recent approaches tackle interference through parameter drop, redundancy reduction, and test-time adjustments.
-For instance, Ties-Merging [63] addresses redundant parameters and sign conflicts; DARE [66] randomly removes redundancies and rescales key parameters; EMR-Merging [18] uses masks and rescaling to retain critical values; AdaMerging [65] learns the importance of task vectors at test time.
-To preserve more task-specific details, WEMoE [50] and Twin-Merging [32] utilize trained routers to adjust shared and task-specific components, while Representation Surgery [64] refines the representation space with test-time training on unlabeled data.
-Overall, existing model merging methods assume shared features alone can capture all task-specific characteristics, but this often fails, losing critical task information.
-Our proposed method addresses this by focusing on task-specific layers rather than just shared representations.
-By refining these layers while keeping the encoder shared, Model Tinting captures essential task-specific features often overlooked in shared layers.
-This approach optimizes merging coefficients and task-specific layers, minimizing conflicts and often surpassing individual model performance.
-Figure 1: Performance comparison of model merging methods. Radar charts highlight Model Tinting’s superiority across (a) 20 vision tasks and (b) 8 NLP tasks, achieving state-of-the-art results.
-3 Preliminary
-3.1 Background
+Negative transfer remains a fundamental challenge in model merging for MTL, as merging often inherits task interference within the shared representation.
+The goal is to construct a single model that generalizes well across all tasks while minimizing such interference.
+Existing weight alignment methods [71, 55, 1] enable merging fine-tuned weights obtained from different pre-trained models.
+However, these approaches face scalability limitations as the number of tasks increases.
+A foundational work in model merging for MTL is Task Arithmetic [23], which introduces task vectors, defined as the difference between pre-trained and fine-tuned models.
+This enables arithmetic-based weight combinations across multiple tasks, allowing the merging of a large number of individual models.
+Subsequent methods refine this approach by explicitly addressing task conflict within the merged model.
+Ties-Merging [72] mitigates redundancy and sign conflicts, while DARE [75] removes redundant components and rescales key parameters.
+AdaMerging [74] learns task vector importance at test time, and LiNeS [64] reduces the impact of shallow layers to preserve general representations.
+While merged models have traditionally been considered task-agnostic, recent findings [63, 19] suggest that they inherently encode task-specific knowledge.
+Motivated by these observations, recent approaches incorporate task-specific modules into the merged encoder to better align with the performance of individual models.
+Surgery [73] refines merged encoders by introducing task-specific adapters, while EMR-Merging [21] applies task-specific masking to a pre-defined merged model.
+Router-based methods [58, 36] determine task vector coefficients via routing mechanisms that depend on each task’s input.
+However, these methods introduce additional parameters and often overlook the role of a well-structured shared expert in maximizing task-specific module effectiveness.
+Model Tinting explicitly refines both shared and task-specific layers in a complementary manner, which allows it to easily capture hidden task-specific knowledge.
+This balanced design improves task adaptation and often surpasses individual model performance.
+3 Background
+3.1 Preliminary
 Problem definition.
@@
 For each of the $K$ tasks, $\hat{z}^{k}$ is fed into a task-specific layer $g(\cdot\ ;\theta_{k})$, where $\theta_{k}$ represents weights unique to task $k$, to perform classification (e.g., the classification head in CLIP [48]).
-Each task output should align with the task nature, and an improved MTL model should provide higher accuracies across tasks.
 We employ two types of weights to address MTL: $\phi_{\text{pre}}$ and $\phi_{k}$, which denote a pre-trained weight and task-specific fine-tuned weights for task $k$, respectively.
@@
 With the merged weight and individual classifiers, a single merged model $f(x;\phi_{\text{MTL}})$ is capable of managing all tasks with decent performance.
-AdaMerging [65] improved the Task Arithmetic by building a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task-wise and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$.
-Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing the objective: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for a test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient.
-Although AdaMerging effectively merges multi-task weights, we believe there is still room for improvement.
+AdaMerging [74] improves Task Arithmetic by constructing a merged model $f(x;\phi_{\text{MTL}})$ through learning $\lambda$ in a task- and layer-wise manner, defined as $\phi_{\text{MTL}}^{l}=\phi_{\text{pre}}^{l}+\sum_{k=1}^{K}\lambda_{k}^{l}\cdot\tau_{k}^{l}$.
+Each merging coefficient $\{\lambda_{k}^{1},\ldots,\lambda_{k}^{L}\}$ is trained by optimizing: $\underset{\{\lambda_{k}^{l}\}}{\min}\sum_{k,x^{k}\in\mathcal{X}}\mathcal{L}_{k}\left(g_{k}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}\right)\right).$ Without ground-truth labels (for test set $\mathcal{X}$), AdaMerging adopts an entropy minimization objective $\mathcal{L}_{k}$ to learn each coefficient.
+Figure 2: Target and cross-task evaluations across 8 tasks using ViT-B/32 with base and trained classifiers.
+We also evaluate both Task Arithmetic and AdaMerging encoders.
+Training classifiers on merged representations improves cross-task generalization while preserving target task performance.
+Additionally, performance varies depending on the shared encoder’s ability to retain task-specific features.
 3.2 Pilot Study
-The success of prior MTL merging methods [32, 66, 63, 65, 18] lies in reducing conflicts in the encoder’s representations across tasks.
-However, we argue that adopting the shared representations without task-specific knowledge may harm MTL performance due to the task-agnostic nature of the merged encoder [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the merged encoder becomes more likely to be task-agnostic..
+The success of recent MTL merging methods [36, 21, 58, 73] lies in introducing additional task-specific components to extract hidden task-specific knowledge [19, 63] from a predefined merged encoder (i.e., shared encoder) [^1] Those MTL merging methods generally achieve overall improvements across tasks through weight averaging, which suggests that the shared encoder becomes more likely to be task-agnostic..
+Instead, we argue that selecting appropriate layers within the model is sufficient without introducing additional modules.
+Furthermore, we suggest that it is crucial to adjust the shared encoder to better capture task-specific information.
 Study outline.
-Our goal is to validate the claim and take it a step further by showing that the following solution works: a trained task-specific layer in a merged model could enhance task-specific capabilities in task-agnostic representation.
-Here, we use the classifier as the layer for analysis, as it is the most straightforward option.
-If our claim is correct, the merging methods [65, 64] could benefit performance-wise from minimal additional training.
-Cross-task evaluation.
-Given that each $k$-th task has its own fine-tuned model, $g_{k}^{bs}\left(f(x^{k};\phi_{k});\theta_{k}^{bs}\right)$, which serves as our control group, where $g_{k}^{bs}(\cdot;\theta_{k}^{bs})$ denotes the base (or primitive) classifier in the pre-trained model (e.g., CLIP [43]).
-We define the cross-task performance, computed by using the visual encoder from another task $m$ in the form $g_{k}^{bs}\left(f(x^{k};\phi_{m});\theta_{k}^{bs}\right)$ for a given task $k$ (i.e., swapping the trained layer only with the original one).
-We emulate the optimization-based merging [65] to observe whether the trained weight would steer towards a task-specific direction.
-To this end, a weight-averaged encoder [19], $f(x^{k};\phi_{\text{MTL}})$ is first employed as the backbone to generate task-agnostic representations.
-Training is performed to minimize the following objective with respect to the designated layers ($g_{k}^{tr}$)’s weight ($\theta_{k}^{tr}$):
+Our goal is to validate this claim and extend it further by demonstrating that a trained task-specific layer can enhance task-specific capabilities even within a task-agnostic representation.
+Additionally, we investigate whether the shared encoder affects the extraction of task-specific knowledge.
+We use the classifier as our task-specific layer for analysis due to its simplicity.
+If our hypothesis holds, merging methods [74, 73] could achieve improved performance with minimal additional training.
+Task-specific training.
+We generate task-agnostic representations from the fixed merged encoder $f(x^{k};\phi_{\text{MTL}})$, which is obtained using Task Arithmetic [23] or AdaMerging [74].
+Then, we train a base classifier $g_{k}(\cdot;\theta_{k})$ from the pre-trained model for the corresponding $k^{th}$ task (e.g., $k=1,...,8$).
+The following objective is minimized with respect to the weights ($\theta_{k}^{tr}$) of the selected layer ($g_{k}^{tr}$) during training:
 $\underset{\{\theta_{k}^{tr}\}}{\min}\sum_{k,(x^{k},y^{k})\in\mathcal{X}}\mathcal{L}_{k}(g_{k}^{tr}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{tr}),y^{k}),$ (1)
-where $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$.
+where $g_{k}^{tr}$ represents the trainable classifier and $y^{k}$ denotes the corresponding ground-truth label for the input sample $x^{k}$.
 This setup uses labeled samples to confirm the encoder’s maximal capability by utilizing ground-truth labels.
-We then compute cross-task scores by pairing each trained classifier $g_{k}^{tr}(\cdot;\theta_{k}^{tr})$ with individual encoders, which is formulated as $g_{k}^{tr}(f_{m}(x^{k};\phi_{m});\theta_{k}^{tr})$ for all $m\neq k$.
-(a) Primitive classifiers
-(b) Trained classifiers
-Figure 2: Cross-task evaluations across diverse encoders and heads from different tasks.
-Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier.
-The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal.
-(a) Primitive (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders.
-(b) Our trained classifiers show mostly reduced performance loss (i.e., higher relative numbers).
-This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise.
-Our interpretation.
-Figure 2 shows cross-task evaluations via individual visual encoders with their respective primitive classifiers; Figure 2 shows results by combining the individual visual encoders with the trained classifiers.
+Target & cross-task evaluation.
+We evaluate each base $g_{k}^{bs}$ and trained classifier $g_{k}^{tr}$ using features from all individual encoders $f(\cdot;\phi_{\text{m}})$ (e.g., $648^{2}=$ 64 cases).
+To assess its effectiveness, we compute the average performance in two settings: 1) target task evaluation, where the classifier is tested on features from its corresponding encoder (i.e., $m=k$), and 2) cross-task evaluation, where it is tested on features from individual encoders trained on distinct tasks (i.e., $m\neq{}k$).
+Interpretation.
+Figure 2 shows the average performance of target task and cross-task evaluations separately, using base classifiers and trained classifiers.
 We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks.
-For instance, integrating the MNIST classifier with the Cars encoder achieves 58.1% accuracy (Figure 2), surpassing the base classifier’s 51.8% (Figure 2).
-These results suggest that training the classifier on merged features for the MNIST task also facilitates the alignment of individual Cars features with the MNIST task.
+Detailed performance for each task pair is provided in Appendix C.1.
+These results suggest that training the classifier on merged features also facilitates the alignment of individual features with other tasks.
 This demonstrates the trained layer’s enhanced capability to refine task-specific outputs from task-agnostic representations.
-We further conjecture that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
-A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels for queries in practice.
-Therefore, our method is designed with this constraint in mind so that it can work in real-world scenarios.
+We further argue that an intermediate layer in the backbone could function similarly, and this will also be empirically proven by our experiments (see Figure 4).
+Additionally, the choice of the shared encoder used for classifier training also affects performance.
+Interestingly, we observe that when using the Task Arithmetic features, cross-task performance improves significantly more than with the AdaMerging features.
+This highlights the importance of identifying a suitable shared encoder that better leverages task-specific knowledge.
+A remaining challenge is that weight merging is typically unavailable to access the ground-truth labels in practice.
+Therefore, our method is designed to handle this constraint and remains effective in real-world scenarios.
 4 Method
-This section introduces our method, inspired by the pilot study, which incorporates task-specific adaptation during weight merging to improve MTL performance significantly.
 4.1 Model Tinting
 Motivation.
-Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific capabilities, potentially boosting overall MTL performance.
-Motivated by this finding, we explore an effective merging method that selectively adapts a few layers in the merged encoder to each task during the merging process.
-We conjecture that this would merit improved MTL performance by injecting task-specific knowledge into the task-agnostic encoder, barely incurring extra training costs.
-Leveraging task-specific knowledge at test-time.
-We design our method based on test-time approaches [65, 64], where ground-truth labels are unavailable.
-A straightforward approach would be to adopt the entropy [65] as the metric replacing any supervisory signals.
-Since the entropy roughly follows the cross-entropy’s trend [65, 38], we believe that the self-entropy may not fully align when data cardinality is insufficient.
-We empirically found that the entropy does not closely approximate the cross-entropy (with ground-truths) during training (see Table 1), so we identify an effective way of employing the cross-entropy.
-Drawing from what we found in the pilot study, we mimic each task encoder’s knowledge to equip the classifier with task-specific knowledge.
-We define an objective to align the distribution of each fine-tuned individual model, $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}\left(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}}\right)|{x^{k}{\in}\mathcal{X}}\}$, with the merged model’s, $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}\left(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}}\right)|{x^{k}{\in}\mathcal{X}}\}$ on test data $\mathcal{X}$.
-While some semi-supervised techniques [68, 51, 41, 1] could be employed, we opt for a simpler approach [29]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach.
+Our pilot study revealed that training a single layer (e.g., a classifier head) on merged representations enhances task-specific performance and can improve overall MTL effectiveness.
+Building on this finding, we propose an efficient merging method that fine-tunes a selected single layer in the merged encoder for each task, extending beyond standard adaptation approaches, while also refining the merged model.
+Task-specific adaptation at test time.
+Our method follows test-time adaptation strategies [74, 73], where ground-truth labels are unavailable.
+A common approach is to use entropy [74] as a proxy for supervision.
+However, entropy only roughly follows the trend of cross-entropy [74, 44] and may not align well when data cardinality is insufficient.
+Our empirical analysis (see Table 2) shows that entropy does not consistently approximate cross-entropy with ground-truth supervision during training.
+Thus, we propose a more effective way to incorporate cross-entropy.
+Building on our pilot study (refer to Section 3.2), we align the classifier’s output with each task encoder’s knowledge to inject task-specific information.
+We define an objective to align each fine-tuned model’s distribution with that of the merged model on the test set $\mathcal{X}$.
+Formally, for each task $k$, we minimize the discrepancy between $C_{k}^{\text{ft}}=\{g_{k}^{\text{bs}}(f(x^{k};\phi_{k});\theta_{k}^{\text{bs}})\mid x^{k}\in\mathcal{X}\}$ and $C_{k}^{\text{merged}}=\{g_{k}^{\text{tr}}(f(x^{k};\phi_{\text{MTL}});\theta_{k}^{\text{tr}})\mid x^{k}\in\mathcal{X}\}$.
+Instead of relying on semi-supervised techniques [77, 59, 46, 2], we adopt a simpler approach [33]: training the merged model to match the fixed output distributions of the individual models [^2] While individual models could be updated during training or replaced with other merged models as alternative options, we opt for a simpler approach.
 We provide the related study in Figure 5.
@@
 $\underset{\{\lambda_{k}^{l}\},\{\theta_{k}^{\text{tr}}\}}{\min}\;\sum_{k=1}^{K}\mathcal{L}_{k}\left(C_{k}^{\text{merged}},C_{k}^{\text{ft}}\right),$ (2)
-where $\mathcal{L}_{k}$ denotes the loss for each task, with $g_{k}^{\text{bs}}$ representing the primitive task-specific layers, and $g_{k}^{\text{tr}}$ representing the learnable task-specific layers.
-Notably, in contrast to AdaMerging [65], our approach jointly optimizes both the merging coefficients and the task-specific layers rather than focusing solely on the merging coefficients.
+where $\mathcal{L}_{k}$ represents the loss for each task, $g_{k}^{\text{bs}}$ are the base task-specific layers, and $g_{k}^{\text{tr}}$ are the learnable task-specific layers.
+Unlike AdaMerging [74], which optimizes only merging coefficients in the shared encoder, our approach jointly optimizes both coefficients and task-specific layers to better tailor representations for distinct tasks.
 SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD
@@
 After 0.32 -0.47 0.79 0.94 0.83 0.41 0.96 0.47
-$\Delta$0.34 -1.19 -0.01 +0.08 -0.09 -0.51 -0.04 -0.09
+$\Delta$ (-0.34) (-1.19) (-0.01) (+0.08) (-0.09) (-0.51) (-0.04) (-0.09)
 Ours Before 0.80 0.86 0.97 1.00 0.98 0.99 1.00 0.84
 After 0.82 0.88 0.99 1.00 0.99 1.00 1.00 0.88
-$\Delta$ +0.02 +0.02 +0.02 +0.00 +0.01 +0.01 +0.00 +0.04
+$\Delta$ (+0.02) (+0.02) (+0.02) (+0.00) (+0.01) (+0.01) (+0.00) (+0.04)
 Table 2: Spearman correlation of losses with ground truth cross-entropy loss.
@@
 We evaluate two losses: ‘Entropy’ (the entropy minimization objective) and ‘Ours’ (our self-label-based cross-entropy loss).
+Loss correlation.
+To validate our approach, we analyze how well different loss functions correlate with ground-truth-based cross-entropy loss, which serves as a strong supervision signal when labels are available.
+We compute the Spearman correlation [^3] Spearman correlation quantifies monotonic relationships and is often used for ranked data or non-linear associations.
+between model-generated losses and the ground-truth-based cross-entropy loss.
+A higher correlation indicates that the loss function provides a reliable training signal, even in the absence of explicit labels.
+Table 2 shows that entropy initially correlates moderately with supervised cross-entropy.
+However, after training, this correlation significantly weakens, and in some cases, entropy diverges sharply (e.g.
+Cars: -1.19 drop).
+This suggests that entropy is not a stable proxy for supervision and may fail to maintain consistency as training progresses.
+This finding aligns with prior studies [74, 44], which observed that entropy-based losses can misrepresent model confidence, leading to unreliable optimization.
+In contrast, our model-generated cross-entropy loss starts with a higher correlation and maintains strong alignment with supervised cross-entropy throughout training.
+Unlike entropy, which fluctuates unpredictably, our approach preserves a stable and reliable training signal across tasks.
+4.2 Model Tinting++
+Refining Model Tinting .
+Model-generated supervisory signals contribute to performance improvement, but not all signals are equally beneficial.
+In particular, cases where the merged model correctly classifies a sample while individual models fail may introduce unreliable training data.
+To address this, we assume that only high-confidence samples from individual models should be used for training.
+Prior studies [16] suggest that well-calibrated models mitigate overconfidence and achieve greater robustness.
+Following this principle, confidence-aware filtering is expected to enhance the stability of the merged model [44].
+Directly setting a fixed confidence threshold is challenging and may introduce unwanted biases.
+Instead, we propose a relative confidence-based filtering strategy: We compare the top-1 probabilities of individual and merged models and exclude samples where the merged model assigns a higher probability than the individual models.
+This ensures that the merged model receives supervision primarily from reliable samples, leading to more stable adaptation.
+As discussed in [43], filtering uncertain cases enhances model robustness, which is expected to improve MTL performance across various tasks.
+(a) Prediction discrepancy
+(b) Overall difference
+Figure 3: Prediction discrepancies between the merged model and individual model.
+(a) The upper bars indicate predictions correctly classified by individual models but misclassified by the merged model; the lower (hatched) bars indicate the opposite.
+A lower upper portion and a higher lower portion indicate a positive impact on the merged model performance.
+(b) represents the overall difference between the upper and lower bars.
+Prediction discrepancy analysis.
+To evaluate the impact of our approach, we analyze discrepancies between merged model predictions and those of individual models, which often serve as a performance upper bound in MTL.
+Figure 3 presents two sets of bar graphs illustrating prediction inconsistencies.
+The upper bars represent samples correctly classified by individual models but misclassified by the merged model, while the lower hatched bars indicate cases where individual models fail but the merged model succeeds.
+A method with lower upper bars and higher lower bars is preferable, as it signifies reduced merging-induced errors and improved generalization.
+Among the evaluated approaches, our method achieves the smallest prediction discrepancy, significantly reducing the number of misclassified samples.
+This suggests that the filtering strategy effectively retains the strengths of individual models while mitigating degradation caused by merging.
+Figure 3 further summarizes the overall impact by computing the net difference between the upper and lower bars.
+A smaller value in this metric indicates a more balanced model that preserves individual model accuracy while capturing additional gains.
+The proposed method consistently outperforms alternatives, demonstrating its effectiveness in minimizing errors while improving generalization, leading to enhanced MTL performance.
 Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg.
 Individual 0.6018 0.9404 0.8922 0.9063 0.9141 0.8720 0.9271 0.7906 0.8556
+[…]
 Weight Averaging 0.1396 0.6411 0.6936 0.3184 0.7536 0.4219 0.5870 0.5523 0.5134
@@
 Task Arithmetic [23] 0.1878 0.8589 0.7990 0.7403 0.8378 0.5908 0.6967 0.6209 0.6665
+[…]
 Ties-Merging [72] 0.2048 0.8440 0.8113 0.5819 0.8570 0.6465 0.7481 0.4296 0.6404
+[…]
 EMR-Merging [21] 0.3996 0.9335 0.8627 0.8277 0.8972 0.8545 0.8957 0.7437 0.8018
@@
 $\dagger$ denotes our reproduced results for a fair comparison.
-Loss correlation.
-Here, we justify our design choice as represented in the loss in Eq.
-(2).
-Inspired by the analysis [65], we demonstrate the effectiveness of using cross-entropy with model-generated labels by computing the Spearman correlation between cross-entropy losses based on individual model signals and ground-truths; the correlation of AdaMerging’s entropy loss versus the ground-truth-based cross-entropy loss is also computed.
-In addition, we further examine the correlation before and after training to see how training affects the correlation.
-Table 1 reports the correlation coefficients for each task.
-We use the weight-averaged model [19] to provide baseline values in the ‘before’ rows, while ‘after’ denotes the correlations measured using a trained merged model.
-We observe that the entropy loss initially shares the correlations with the supervised cross-entropy loss, and the correlation mostly dropped after training across diverse tasks (i.e., 1 out of 8), particularly on challenging ones like Cars, even showed a negative correlation.
-This suggests that entropy minimization may not fully follow the true supervisory signals for MTL [^3] As also observed in prior studies [65, 38], entropy sometimes diverges further from the supervised cross-entropy for certain sample distributions.; the resulting merged model does not appear to be trained under ground-truth supervisory signals.
-In contrast, our model-generated cross-entropy loss initially shows higher correlations and maintains a striking, strong positive correlation after training.
-Intriguingly, the correlation strengthens after training, which suggests that the merged model becomes more aligned with supervisory signals, likely due to its improvement under appropriately emulated signals.
-Therefore, our approach effectively trains a merged model that can better follow the ground-truth signals at test time when labels are inaccessible.
-4.2 Model Tinting++
-Improving Model Tinting.
-Model-generated supervisory signals are crucial for our method, and enhancing these signals could yield further improvements.
-Our insight is to exclude undesirable samples during training where the merged model is correct while individual models are incorrect in that training with such signals may not always be beneficial.
-How many such samples exist, and what is their impact?
-Figure 5 shows the extent to which the number of such cases is not negligible.
-Another question could be: how can we identify these samples for exclusion?
-First, we presume only confident samples for individual models would be meaningful in training.
-Additionally, we refer to the literature [15], which suggests that a calibrated model achieves robustness by mitigating overconfidence; we similarly assume that a merged model would follow the trend and less-confident samples would be helpful [38].
-Since quantifying confidence is challenging and introducing a new confidence threshold is undesirable, we sample based on the confidence trends in both individual and merged models.
-Specifically, we use the top-1 probability as the metric and filter out the instances when the merged model’s top-1 probability for a given prediction is higher than that of the individual model.
-This approach can be understood as selecting only the samples that could calibrate the merged model while individual models provide reliable supervisory signals for training.
-This ensures that the resultant merged model becomes more robust, as similarly discussed in [39]; thereby, we expect improved MTL performance across different tasks.
-Our experimental results will demonstrate the improvements achieved by the proposed method.
-Dicussions on implementation.
-Previous methods [65, 50] sample a batch for each task, pass each batch through a shared backbone, and then process it through the corresponding task-specific head, which is typical.
-The losses from each task are then combined, and a single gradient update is performed after completing the forward pass for all tasks.
-However, this approach becomes less efficient as the number of tasks or model size increases.
-To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one.
-Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach.
-While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6.
-Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%).
-We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [26] by training tasks sequentially rather than simultaneously.
 Method SUN Cars RES. Euro SVHN GTS. MNIST DTD Avg.
@@
 Ties Merging [72] 65.0 64.4 74.8 77.4 81.2 69.3 96.5 54.5 72.9
+[…]
 […]
@@
 […]
+[…]
 Concrete TA [56] 74.6 86.2 89.0 96.7 93.6 93.4 99.1 66.9 87.4
@@
 (b) On merging ViT-L/14 models
-Table 3: Multi-task performance comparison across 8 vision tasks. We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models.
+Table 4: Multi-task performance comparison across 8 vision tasks.
+We compare our methods with the competing methods for merging ViT-B/32 (left) and ViT-L/14 (right) fine-tuned models.
+$\dagger$ denotes reproduced results.
 5 Experiment
@@
 Datasets and metrics.
-We follow the standard setups in DARE [66] and EMR-Merging [18] for NLP experiments: employing fine-tuned weights for 8 tasks from the GLUE benchmark [54].
-Evaluation on each task is performed with task-specific metrics: Matthews correlation for CoLA [56], Pearson and Spearman correlations for STS-B [4], and accuracy for SST-2 [47], MRPC [11], QQP [20], MNLI [57], QNLI [44], and RTE [13].
-For vision tasks, we follow the standard Task Arithmetic [19] and AdaMerging [65] setups, using fine-tuned weights on 8 standard image classification tasks.
-Furthermore, we extend our vision task experiments with 12 more tasks reaching the overall experimental scope to 20 tasks, following [55]: CIFAR-10 [28], CIFAR-100 [28], FER2013 [14], Flowers102 (Flowers) [37], Oxford-IIIT Pet (Pet) [40], PCAM [53], STL10 [8], EMNIST [9], FashionMNIST (FMNIST) [60], Food101[2], KMNIST [7], and Rendered SST-2 (R-SST) [42].
-All tasks use the accuracy metric.
-Additionally, we provide results for a selected subset of 14 out of the 20 tasks in Appendix.
+We follow the standard setups in DARE [75] and EMR-Merging [21] for NLP tasks, using fine-tuned weights for eight tasks from the GLUE benchmark [62].
+Task-specific metrics are employed: Matthews correlation for CoLA [65], Pearson and Spearman correlations for STS-B [5], and accuracy for other tasks [53, 12, 24, 66, 49, 14].
+For vision tasks, we adopt the setups from Task Arithmetic [23] and AdaMerging [74], using fine-tuned weights on eight image classification tasks.
+We extend our experiments to 12 additional tasks [32, 15, 42, 45, 61, 9, 10, 69, 3, 8, 47], covering a total of 20 tasks as in [63].
+All vision tasks are evaluated using accuracy.
+Results for a subset of 14 tasks are provided in the Appendix C.3.
+For dense prediction evaluation, we use NYUv2 [52], which contains 1,449 RGB-D indoor scene images.
+It includes three tasks: 13-class semantic segmentation, depth estimation, and surface normal estimation.
+We evaluate them with mIoU/pixel accuracy, absolute/relative error, and mean angular error, respectively.
 Models.
-We employ RoBERTa-base [31] for the NLP tasks in our experiments.
-We use the publicly released weights [^4] https://github.com/harveyhuang18/EMR_Merging by the prior work [18].
-For a fair comparison, we run experiments for TWIN-Merging [32] and our method using fine-tuned weights from EMR-Merging.
-Our method applies Mean Squared Error (MSE) loss to STS-B instead.
-For vision experiments, we utilize the pre-trained CLIP [43] weights [^5] https://github.com/mlfoundations/open_clip, specifically ViT-{B/32, B/16, L/14}.
-Most analyses and ablation studies are performed with ViT-B/32 for comparison.
-For a fair comparison across the 8 tasks, we use the fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors publicly released by Task Arithmetic [19].
-In experiments with 14 and 20 vision tasks, CLIP ViT-B/32 is fine-tuned based on the configuration used in Task Arithmetic to obtain individual models.
-Due to the lack of 14 and 20 tasks experiments in literature, we reproduce Task Arithmetic, MagMax [34], Ties Merging [63], Consensus TA [55], AdaMerging [65], Surgery [64], and EMR-Merging [18].
-More detailed experimental setups are provided in Appendix.
-5.2 Main Results
-Merging 8 NLP tasks.
-As shown in Table 2, our Model Tinting largely improved over the best method on average.
-While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks.
-We argue that this reveals the limitations of methods that only handle shared representations and shows the superiority of our method that effectively manages task-specific information.
-For further details, please refer to Appendix.
-Merging 8 vision tasks.
-Table 3 and Table 3 demonstrate that our approach surpasses traditional MTL methods and achieves average performance improvements of 0.7%p and 0.2%p on ViT-B/32 and ViT-L/14, respectively, compared to the previous state-of-the-art merging methods.
-(For ViT-B/16, we observe an improvement of 0.7%p; more details are provided in Appendix) Additionally, our method achieves performance close to the individual models, even surpassing them on some tasks.
-When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal.
-This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
-Merging 20 vision tasks.
-In Table 4, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance.
-Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic.
-In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p.
-Notably, for several tasks, our method even exceeds the performance of individual models.
-With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
+For NLP tasks, we use RoBERTa-base [35], leveraging publicly available fine-tuned weights [^4] https://github.com/harveyhuang18/EMR_Merging from EMR-Merging [21].
+We conduct experiments on TWIN-Merging [36] and our method under the same settings, using Mean Squared Error (MSE) loss for STS-B.
+For vision experiments, we employ pre-trained CLIP [48] models [^5] https://github.com/mlfoundations/open_clip, specifically ViT-B/32, B/16, L/14.
+Most analyses use ViT-B/32.
+To ensure comparability across eight tasks, we use publicly available fine-tuned weights [^6] https://github.com/mlfoundations/task_vectors from Task Arithmetic [23].
+For 14- and 20-task experiments, we fine-tune CLIP ViT-B/32 following the Task Arithmetic configuration.
+For dense prediction, we adopt ResNet-50 [18] as the backbone, following FusionBench [57].
+The model is initialized with ImageNet pre-trained weights and fine-tuned for each task.
+Since prior works lack results on 14- and 20-task settings, we reproduce multiple previous methods [38, 72, 64, 63, 74, 73, 21].
 Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD CIFAR100 FER2013
@@
 Task $\dagger$ [23] 61.8 53.0 61.9 57.5 49.8 44.6 77.9 45.6 65.4 41.4
+[…]
 […]
@@
 […]
+[…]
 Model Tinting
@@
 […]
+[…]
 Model Tinting
@@
 Table 5: Task-specific and average performance across 20 vision tasks. We use ViT-B/32 for experiments. $\dagger$ denotes reproduced results.
+Method Segmentation Depth Estimation Normal
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Model Tinting
+49.0 73.0 45.7 18.6 26.4
+Model Tinting++
+50.2 73.4 45.1 18.6 26.1
+Table 6: Multi-task performance across three dense prediction tasks.
+We compare the performance of merging ResNet50 models fine-tuned on three dense prediction tasks in NYUv2.
+$\dagger$ denotes reproduced results.
+5.2 Main Results
+Merging 8 NLP tasks.
+As shown in Table 3, our Model Tinting largely improved over the best method on average.
+While EMR-Merging suffers from a significant performance drop on CoLA, and TWIN-Merging on STS-B, Model Tinting demonstrates consistent performance close to individual task-specific models across all tasks.
+We argue that this reveals the limitations of methods that rely on a fixed shared encoder and shows the superiority of our method, which effectively manages task-specific information from the adaptive shared encoder.
+Merging 8 vision tasks.
+Table 4 and Table 4 show that our method outperforms traditional MTL methods, improving ViT-B/32 and ViT-L/14 by 0.7%p and 0.2%p over state-of-the-art methods (ViT-B/16 improves by 0.7%p; see Appendix C.3).
+Notably, for ViT-B/32, this is achieved with just 0.5M trainable parameters, far fewer than the 7M used in prior state-of-the-art approach.
+Additionally, our method achieves performance close to the individual models regardless of task difficulty and even surpasses them on certain tasks.
+For instance, on DTD, a particularly challenging dataset, other merging methods exhibit a noticeable performance gap compared to individual models, whereas our approach achieves a higher accuracy of 80.2% compared to 79.4%.
+When applying sample filtering, we observe a 0.5%p improvement for ViT-B/32, while the performance gain for ViT-L/14 is minimal.
+This is likely because Model Tinting already performs nearly as well as individual models when using ViT-L/14.
+Merging 20 vision tasks.
+We explore the impact of scaling model merging by expanding the task set from 8 to 20, making the problem more complex.
+To ensure a fair comparison, we reproduce the performance of existing methods under the same experimental conditions.
+In Table 5, we observe that for KMNIST, both task arithmetic and AdaMerging exhibit notably low performance.
+Furthermore, on datasets like PCAM and EMNIST, AdaMerging occasionally results in lower performance than task arithmetic.
+In contrast, our approach shows a significant improvement, with a 19.3%p performance increase over AdaMerging and surpassing state-of-the-art methods by 2.3%p.
+Notably, for several tasks, our method even exceeds the performance of individual models.
+With sample selection, we see an additional 0.5%p performance boost, highlighting the effectiveness of our approach as the number of tasks scales.
+Merging dense prediction tasks.
+Dense prediction tasks, such as segmentation, depth estimation, and normal estimation, pose significant challenges due to their inter-task heterogeneity.
+Despite this, most model merging studies focus on classification, leaving these tasks largely unexplored.
+Table 6 presents results for these tasks, highlighting the limitations of previous methods.
+Surgery, which performs comparably to individual models on eight classification tasks, suffers substantial performance drops in depth and normal estimation.
+Similarly, EMR-Merging, despite leveraging task-specific dynamic masks, struggles to maintain segmentation performance.
+In contrast, our approach effectively balances these tasks, maintaining performance close to individual models even in dense prediction settings.
 5.3 Emprical Analyses
+Effect of adjusting different layers.
+A natural question is whether the improvements observed in Model Tinting are specific to training the classifier.
+To explore this, we analyze the effect of training different layers beyond the classifier by updating merging coefficients along with each of the 12 transformer layers in the merged encoder.
+As shown in Figure 4, training a single internal layer achieves accuracy comparable to training the classifier, suggesting that task-specific information can be effectively captured by refining a single layer rather than modifying the entire model.
+Identifying the optimal layer could further enhance performance.
+On the other hand, training multiple layers at once, either from the early stage (layers 1–6) or the later stage (layers 7–12), leads to a significant drop in performance.
+This suggests that updating too many layers disrupts the task-agnostic knowledge embedded in the pre-trained model, reducing its ability to generalize across tasks.
+Figure 4: Impact of training different task-specific layer.
+We compare the average performance of merged models when different task-specific layers are trainable, while other layers update only their merging coefficients.
+Training a single layer achieves accuracy similar to tuning the classifier, whereas modifying multiple layers leads to performance degradation.
 Sparsity analysis of merging coefficients.
-Inspired by the literature [66, 63, 10] in which redundant parameters degrade performance due to task-specific parameters’ conflicts, we analyze the learned merging coefficients to observe this effect in our method.
-We explore sparsity in the merging coefficients after training them alone and jointly with the classifier.
-Figure 3 shows both configurations sequentially and demonstrates that the joint training enjoys striking sparsity.
+Inspired by the literature [75, 72, 11], which highlights how redundant parameters degrade performance due to conflicts among task-specific parameters, we investigate this phenomenon in our method.
+We analyze the learned merging coefficients to observe their impact, specifically exploring sparsity after training them alone and jointly with the classifier.
 We find 37.2% versus 55.9% of the coefficients are concentrated near zero (i.e., smaller than 1e-5) with the deviated average accuracy of 83.4% versus 89.6% for each configuration.
 This suggests that training a task-specific layer, like the classifier, along with merging coefficients, reduces task interference by pruning unnecessary parameters, leading to improved accuracy.
-(a) Training coefficients only
-(b) Joint training
-Figure 3: Impact of joint training on coefficient sparsity.
-We visualize the merging coefficients to display the sparsity in weights.
-Joint training of merging coefficients with a task-specific layer promotes sparsity strikingly, with 55.9% of coefficients near zero compared to 37.2% when training coefficients only.
-Can inner layers serve as task-specific adjustments?
-One might wonder if our observed improvements are unique to training just the classifier.
-We disprove this speculation by evaluating other layers beyond the classifier.
-Specifically, we train each of the 12 transformer layers in a merged encoder, which is performed under a similar setup in the pilot study.
-Figure 4 shows transformer layers generally achieve performance matched to training the classifier (see the details in Appendix).
-This suggests that training a single internal transformer layer can do similarly to what the classifier does.
-Furthermore, identifying the optimal layer could lead to additional performance improvements.
-Meanwhile, when we train either all layers or groups of layers — early layers (from layers 0 to 5-th) or late layers (from layers 6 to 11-th) together — performance drops significantly; we believe that training many layers at once in MTL scenarios diminishes the task-agnostic knowledge inherently stored in a pre-trained model.
-Figure 4: Impact of training different task-specific layers.
-Performance comparison of merged models when training: 1) a single inner layer, 2) late layers, 3) early layers, and 4) only the classifier.
-Training a single layer matches the performance of training only the classifier, while training multiple layers reduces accuracy.
-(a) Prediction discrepancy
-(b) Overall difference
-Figure 5: Prediction discrepancies between the merged model and individual model.
-(a) The upper bars indicate predictions correctly classified by individual models but misclassified by the merged model; the lower (hatched) bars indicate the opposite.
-A lower upper portion and a higher lower portion indicate a positive impact on the merged model performance.
-(b) represents the overall difference between the upper and lower bars.
-Assesing merging methods via prediction discrepancy.
-Here, we assess diverse merged models with their sample prediction correctness with respect to the individual models that often serve as the performance upper bound in MTL due to higher accuracy.
-Figure 5 displays two sets of bar graphs.
-First, the upper bars represent samples correctly predicted by the individual models but incorrectly by the merged model, where a lower one aligns more with the individual models.
-Our method achieves the smallest discrepancy, significantly outperforming AdaMerging and Surgery.
-The lower bars in Figure 5 represent samples misclassified by the individual models but correctly classified by the merged model, representing unexpected gains of the merged model.
-A higher one here benefits from the merged model’s predictions beyond those of the individual model.
-Figure 5 illustrates the overall impact by computing the difference between the penalties and gains shown in the upper and lower bars of Figure 5.
-First, our improved method presented in Section 4.2 successfully enhances this metric, where a lower one is better.
-We believe our methods effectively balance preserving the strengths of the individual models while capturing unexpected gains, leading to superior overall performance.
-We observe that while closely related methods like AdaMerging [65] (using entropy minimization without constraints) and Surgery [64] (performs post-AdaMerging to preserve individual model predictions) also enjoy some unexpected gains, the penalties observed in the upper bars outweigh the benefits, which incur net decreases in the overall metric.
-(a) On supervisory model
-(b) Training robustness
-Figure 6: More analyses with merging coefficients.
-(a) Refined supervisory models merged under various coefficients are tested.
-Our design choice – using the unmerged individual models – performs near-optimally.
-(b) MTL accuracy confirms the initialization robustness of our method across different initial merging coefficients; related results are illustrated in Figure 7.
+For visualizations of merging coefficients, please refer to Appendix D.4.
 More studies with merging coefficients.
@@
 Efficiently searching for the coefficient would be a potential research direction in the future.
+(a) On supervisory model
+(b) Training robustness
+Figure 5: More analyses with merging coefficients.
+(a) Refined supervisory models merged under various coefficients are tested.
+Our design choice – using the unmerged individual models – performs near-optimally.
+(b) MTL accuracy confirms the initialization robustness of our method across different initial merging coefficients; related results are illustrated in Figure 6.
 Initialization robustness.
-Our method benefits from the robustness of the merging coefficients’ initialization.
+Our method benefits from robust merging coefficient initialization.
 To highlight this, we present training curves under different initialization settings and compare them with those of other model merging methods.
-As shown in Figure 6, when only the classifier is trained, the performance is sensitive to initialization; however, when both the classifier and the coefficients are trained together, the method demonstrates substantial robustness.
-Figure 7 shows that when training a Surgery model or the classifier only, starting from the AdaMerging initialization yields higher performance than starting from the Task Arithmetic initialization, which suggests the sensitivity to initialization.
+As shown in Figure 5, when only the classifier is trained (i.e., fixed shared encoder), the performance is sensitive to initialization; However, training both the classifier and coefficients together improves robustness.
+Figure 6 shows that when training a task-specific adapter only, the Surgery model is sensitive to initialization.
+Unlikely, our method jointly optimizes the merging coefficient and task-specific layer, making it robust to initialization sensitivity.
 Figure 6: Impact of initializations.
@@
 6 Conclusion
-In this paper, we have introduced Model Tinting , a novel approach to multi-task model merging that enhances task-specific knowledge through adaptively learning task-specific layers, departing from traditional methods that rely on shared representations.
-Model Tinting++ further involved a novel confidence-based sampling method, which has proven favorable.
-Extensive experiments demonstrated state-of-the-art performance across various domains, including vision and NLP, with particularly significant improvements on the challenging 20-task vision benchmark.
-Our empirical studies further justified that our design choice of incorporating just a single task-wise learned layer is sensible.
-Finally, we discovered additional insights: 1) inner layers all play similar roles in task-specific adaptations; 2) training task-specific layers naturally promotes sparsity in merging coefficients; 3) our joint training is insensitive to initializations with improved results.
-We believe our study could be a foundation for future research in model merging.
+In this paper, we have introduced Model Tinting , an efficient multi-task model merging approach that enhances task-specific knowledge with minimal overhead.
+By adaptively learning task-specific layers while refining the shared model, our method balances generalization and specialization without additional task-specific modules.
+This distinguishes our method from conventional model merging methods, which often introduce unnecessary complexity or struggle to retain fine-grained task information.
+To further improve robustness, Model Tinting++ incorporates a confidence-based filtering mechanism that selectively removes unreliable supervision signals.
+Extensive experiments across vision, NLP, and dense prediction tasks demonstrate that our approach consistently achieves state-of-the-art performance, particularly excelling in dense prediction and large-scale multi-task benchmarks.
+Additionally, our analysis reveals key insights: 1) Inner layers complement each other in capturing task-specific features; 2) Training task-specific layers induces sparsity in merging coefficients, improving task separation; 3) Our joint optimization strategy is robust to initialization, ensuring stable results across tasks.
+We believe this work provides a strong foundation for future research in scalable and adaptive model merging.
 Appendix
@@
   • Appendix B: Detailed experimental setup and hyperparameter configurations;
-  • Appendix C: Additional results for NLP and vision tasks;
+  • Appendix C: Additional results for pilot study, NLP, and vision tasks;
   • Appendix D: Further analyses on task-specific layers, loss functions, convergence, and sparsity.
 Appendix A Overview of Model Tinting
-We give a brief overview of our proposed method Model Tinting , a test-time approach for merging multiple fine-tuned models by introducing a single trainable layer for each task.
-Figure A provides an illustrative overview of our method.
-At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using merging coefficients.
+We give a brief overview of our proposed method Model Tinting , a test-time method that merges multiple fine-tuned models using a single trainable layer per task Figure A provides an illustrative overview of our method.
+At the top, the figure illustrates the merging process, where multiple fine-tuned models ($\phi_{1},\phi_{2},\ldots,\phi_{K}$) are combined into a shared representation ($\phi_{\text{MTL}}$) using adaptive merging coefficients.
 During this process, a trainable task-specific layer ($\phi_{\text{tint}}^{l}$) is introduced at a chosen transformer layer $l$, capturing task-specific information from the task-agnostic representations of the merged encoder.
 We argued that any layer can be picked for task adjustments before merging, and this is further supported by the graph in Figure A.
-The results demonstrate that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline.
+The results show that applying Model Tinting, regardless of the chosen layer, leads to significant performance improvements over the baseline.
 This highlights the robustness and effectiveness of task-specific adjustments in enhancing task-agnostic representations, with more details provided in Appendix D.1.
@@
 For vision tasks, the learning rate is set to 0.01 when training the classifier and 0.001 when training only the merging coefficients.
-For NLP tasks, a learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated.
+For dense prediction tasks, the learning rate of 0.0001 is used.
+For NLP tasks, the learning rate of 0.0005 is used, regardless of whether the classifier, the merging coefficients, or both are being updated.
 Consistent with prior works [74, 58, 73], vision tasks are trained with 16,000 test samples per task over 500 iterations, using a batch size of 32.
-Similarly, NLP tasks are trained with 32,000 samples per task over 1,000 iterations, also with a batch size of 32.
+Dense prediction tasks use 400 samples per task over 25 iterations with a batch size of 16.
+NLP tasks are trained with 32,000 samples per task over 1,000 iterations with a batch size of 32.
 For each task, the loss is computed, and the model parameters are updated immediately following the forward pass for each batch.
 To enable sequential processing, the order of tasks is randomized.
+Figure A: Illustration of   Model Tinting .
+Our method improves multi-task learning by introducing a trainable, task-specific layer during model merging.
+Our approach (red) consistently achieves better results than the baseline without tinting (green) across all layers.
+Dicussions on implementation.
+Previous methods [74, 58] sample a batch for each task, pass each batch through a shared backbone, and then process it through the corresponding task-specific head, which is typical.
+The losses from each task are then combined, and a single gradient update is performed after completing the forward pass for all tasks.
+However, this approach becomes less efficient as the number of tasks or model size increases.
+To address this, we experiment with a sequential update strategy, where updates are applied immediately after each task’s forward pass instead of the typical one.
+Interestingly, this sequential update approach gives a positive byproduct: performance improvements over the traditional approach.
+While the original AdaMerging achieves an average accuracy of 80.1%, our sequential update strategy improves it to 81.6% for the 8 vision tasks using ViT-B/32.
+Similarly, our method also enjoys a gain when using sequential updates (89.6% compared to 89.4%).
+We argue that this improvement may result from mitigating catastrophic forgetting in continual learning [30] by training tasks sequentially rather than simultaneously.
 Pilot study.
-The aforementioned training details are applied almost identically to the experiments in the pilot study presented in our main paper.
-As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic-merged weights for the 8 vision tasks using ViT-B/32.
+The aforementioned training details are applied almost identically to the experiments in the pilot study of our main paper.
+As stated in the main paper, only the classifier was chosen (as a straightforward and practical option) to train on the given Task Arithmetic (or AdaMerging [74]) merged weights for the 8 vision tasks using ViT-B/32.
 A difference is that, instead of using the test dataset, the training dataset is used in a supervised manner to produce the results closer to the upper bound, in which the difference stands out more.
@@
 B.2 Baseline Details
-Figure A: Illustration of   Model Tinting .
-Our method improves multi-task learning by introducing a trainable, task-specific layer during model merging.
-Our approach (red) consistently achieves better results than the baseline without tinting (green) across all layers.
 We compare our method against a diverse set of baselines, ranging from simple merging strategies to advanced methods leveraging task-specific or shared representations.
@@
 It involves removing redundant parameters from the task vector and resolving parameter sign conflicts.
+LiNeS [64] observes that reducing the influence of shallow layers helps prevent distortion of general representations, thereby simplifying merging coefficient selection by allowing them to increase linearly with layer depth.
 Consensus TA [63] enhances Task Arithmetic [23] by using Consensus Merging, which retains only weights beneficial to multiple tasks while eliminating irrelevant or task-specific weights.
@@
 During inference, the router integrates shared and task-specific components dynamically based on input.
+(a) Fixed Classifier
+(b) Classifier Training with Task Arithmetic
+(c) Classifier Training with AdaMerging
+Figure B: Cross-task evaluations across diverse encoders and heads from different tasks.
+Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier.
+The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal.
+(a) Base (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders.
+(b) Classifiers trained with Task Arithmetic features show mostly reduced performance loss (i.e., higher relative numbers).
+This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise.
+(c) Classifiers trained with AdaMerging features also improve over (a) but perform worse than (b), highlighting the importance of a shared encoder when leveraging task-specific knowledge.
+Appendix C More Experimental Results
+C.1 Pilot Study
+Figure 2(a) shows cross-task evaluations via individual visual encoders with their respective base classifiers; Figure 2(b) shows results from combining individual visual encoders with trained classifiers from Task Arithmetic, while Figure 2(c) presents results using the AdaMerging merged encoder.
+We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks.
+For instance, integrating the GTSRB [54] classifier trained on Task Arithmetic features with the SVHN [41] encoder achieves 56.3% accuracy (Figure 2(b)), surpassing the base classifier’s 31.4% (Figure 2(a)).
+These results suggest that training the classifier on merged features for the GTSRB task also facilitates the alignment of individual SVHN features with the GTSRB task.
+This demonstrates the trained layer’s enhanced ability to refine task-specific outputs from task-agnostic representations.
+On the other hand, classifiers trained on AdaMerging features achieve 49.0% accuracy (Figure 2(c)), surpassing the base classifier but performing worse than the classifier trained on Task Arithmetic features.
+This suggests that even when task-specific layers are trained, the choice of a shared encoder can lead to suboptimal performance.
+Therefore, it highlights the importance of the shared encoder in effectively leveraging task-specific knowledge.
 Method SUN Cars RES. Euro SVH. GTS. MNI. DTD C100 FER Flower Pet PCAM STL Avg.
 Pretrained 62.3 59.7 60.7 45.5 31.4 32.6 48.5 43.8 64.2 39.0 66.3 87.4 60.6 97.1 57.1
+[…]
 […]
@@
 […]
+[…]
 AdaMerging++ [74] 71.8 80.8 84.1 94.3 91.9 94.5 98.7 69.8 85.7
@@
 Table B: Multi-task performance when merging ViT-B/16 models on 8 tasks.
-Appendix C More Experimental Results
-Figure B: Multi-task performance when merging RoBERTa models on 8 tasks.
+Figure C: Multi-task performance when merging RoBERTa models on 8 tasks.
 Yellow indicates training only coefficients, green indicates training only the classifier, and blue indicates training both in our method.
-C.1 NLP Tasks
+C.2 NLP Tasks
 We conduct experiments to investigate how the training of the classifier and merging coefficients, respectively, affects the performance of NLP tasks in Model Tinting , as presented in Table 3 of the main paper.
-Figure B shows the results of merging RoBERTa [31] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss.
+Figure C shows the results of merging RoBERTa [35] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss.
 As mentioned in the main paper, evaluation metrics differ across tasks: the Matthews correlation coefficient is used for CoLA [65], the average of Pearson and Spearman correlations is applied to STS-B [5], and the accuracy is used for all other tasks [53, 12, 24, 66, 49, 14].
@@
 Furthermore, training both the classifier and coefficients together demonstrated a complementary effect, achieving the highest performance.
-C.2 Vision Tasks
+C.3 Vision Tasks
 Table A presents the results of merging ViT-B/32 across 14 tasks.
@@
 While other methods exhibit significant performance gaps compared to individual models, particularly on datasets like Cars and DTD, Model Tinting consistently achieves performance close to that of individual models.
-Figure C: Impact of training different task-specific layers with merging coefficients.
-We compare average performance for merged models with different trainable task-specific layers while other layers update only their merging coefficients.
-Appendix D More Empirical Analyses
-D.1 On Task-specific Layers
-Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
-0 66.2 71.9 92.2 99.0 94.2 96.5 99.1 77.0 87.0
-1 65.1 69.9 92.4 98.8 94.2 97.4 99.0 77.6 86.8
-2 64.4 69.8 92.3 99.0 93.8 97.3 98.6 77.4 86.6
-3 64.3 68.3 92.2 99.1 92.8 98.3 99.1 78.2 86.5
-4 64.8 69.5 93.3 98.6 94.6 98.1 99.2 78.2 87.0
-5 64.6 70.7 91.3 99.0 94.8 98.3 99.1 78.8 87.1
-6 67.4 70.7 94.4 98.9 92.3 98.3 99.2 79.1 87.5
-7 67.1 66.8 94.4 99.7 94.8 99.2 99.3 79.2 87.6
-8 66.1 69.8 93.6 99.0 94.2 98.2 99.5 78.6 87.4
-9 68.2 66.5 94.6 99.6 91.9 98.5 99.2 79.5 87.3
-10 68.3 73.0 94.0 99.6 92.5 97.9 99.1 79.4 88.0
-11 68.2 69.0 94.2 99.4 92.3 97.5 99.0 79.3 87.4
-Classifier 69.9 73.8 92.2 98.6 89.5 97.5 99.0 78.8 87.4
-Late 50.7 63.4 89.3 99.6 93.4 95.9 99.3 75.5 83.4
-Early 42.5 54.2 85.0 93.8 93.2 96.3 98.2 76.0 79.9
-All 25.5 31.2 82.7 97.0 92.4 96.2 97.5 72.7 74.4
-(a) Training layers.
 Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
@@
 All 67.1 69.5 94.2 98.4 95.1 97.7 98.8 75.4 87.0
-(b) Training layers with merging coefficients.
-Table C: Performance details of merged ViT-B/32 models when training different layers for task-specific adjustments.
+Table C: Performance details of merged ViT-B/32 models when training different layers with merging coefficients for task-specific adjustments.
 We observe that consistently high performance can be achieved regardless of which task-specific single layer is trained.
-Note that the 0-th layer in the table refers to the first transformer layer.
+Appendix D More Empirical Analyses
+D.1 Detailed Results on Task-specific Layers
 Our proposed method, Model Tinting , introduces the concept of incorporating a task-specific layer to balance shared and task-specific representations during model merging.
 This approach allows effective adaptation while maintaining efficiency to address task conflicts in multi-task learning.
-Building on the results in Figure 4 of the main paper, which demonstrate the potential of training inner layers while keeping the merging coefficients fixed, we extend this analysis further.
-Specifically, we explore whether performance remains robust when both merging coefficients and task-specific layers are trained jointly.
-ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers again.
+ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers.
 When a particular layer is marked as task-specific, it is split into unique versions for each task, resulting in eight task-specific layers.
 While these layers are trained exclusively on corresponding task data, the remaining layers update only the merging coefficients using data from all tasks.
-Figure C shows that training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier.
+Table C presents detailed task-level results.
+‘Early’ refer to layers 1 to 6-th, while ‘Late’ refer to layers 7 to 12-th.
+Training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier.
 These findings highlight the minimal impact of task-specific layer choice on overall performance, reaffirming the robustness of our approach.
-Table C provides detailed task-level results, while Table C shows consistent average performance across layers, with specific layers excelling on certain datasets (e.g., the 11th layer on SUN397 [61] and the 7th layer on GTSRB [48]).
-Conversely, Table C demonstrates that merging coefficients adaptively balance representations across tasks, preventing biases even when other layers remain frozen.
+The results also show consistent average performance across layers, demonstrating that merging coefficients effectively balance representations across tasks, mitigating biases even when other layers remain frozen.
 D.2 On Loss Functions
 Figure D: Comparison of training with various loss functions.
-The yellow bars represent the case where only the coefficients are trained, while the blue bars indicate the joint training of the classifier and coefficients.
+The yellow bars represent training only the coefficients, while the blue bars indicate the joint training of the classifier and coefficients.
 The dashed line corresponds to training with ground truth labels using cross-entropy loss, which is the upper bound.
@@
 After 0.38 0.31 0.94 0.92 0.88 0.96 0.98 0.38
-$\Delta$0.35 -0.51 +0.1 +0.06 -0.08 +0.05 -0.02 -0.23
+$\Delta$ (-0.35) (-0.51) (+0.1) (+0.06) (-0.08) (+0.05) (-0.02) (-0.23)
 Ours Before 0.82 0.92 0.98 1.00 0.99 1.00 1.00 0.85
 After 0.86 0.95 1.00 1.00 0.99 1.00 1.00 0.90
-$\Delta$ +0.04 +0.03 +0.02 +0.00 +0.00 +0.00 +0.00 +0.05
+$\Delta$ (+0.04) (+0.03) (+0.02) (+0.00) (+0.00) (+0.00) (+0.00) (+0.05)
 (a) ViT-B/16
@@
 After 0.63 0.76 0.94 0.81 0.79 0.95 0.82 0.83
-$\Delta$0.22 -0.17 -0.01 -0.17 -0.17 -0.01 -0.18 +0.04
+$\Delta$ (-0.22) (-0.17) (-0.01) (-0.17) (-0.17) (-0.01) (-0.18) (+0.04)
 Ours Before 0.87 0.96 0.98 1.00 0.98 1.00 1.00 0.89
 After 0.89 0.97 1.00 1.00 0.99 1.00 1.00 0.91
-$\Delta$ +0.02 +0.01 +0.02 +0.00 +0.01 +0.00 +0.00 +0.02
+$\Delta$ (+0.02) (+0.01) (+0.02) (+0.00) (+0.01) (+0.00) (+0.00) (+0.02)
 (b) ViT-L/14
@@
 These findings are consistent with the results reported for ViT-B/32 in the main paper, suggesting that our self-labeling approach with the cross-entropy loss could be more effective than the entropy loss when training the classifier alongside the coefficients.
+(a) Comparison of average accuracy for ViT-B/32, ViT-B/16, and ViT-L/14 across training iterations.
+(b) Average accuracy of ViT-B/32 across training iterations under different learning rates.
+Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates. Our method demonstrates robust and fast convergence, regardless of model size and learning rate.
 D.3 Convergence Analysis
@@
 These observations demonstrate the robustness and efficiency of Model Tinting across different model scales and hyperparameter settings.
-(a) Comparison of average accuracy for ViT-B/32, ViT-B/16, and ViT-L/14 across training iterations.
-(b) Average accuracy of ViT-B/32 across training iterations under different learning rates.
-Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates.
-(a) Training coefficients only
-(b) Joint training
-Figure G: Impact of joint training on coefficient sparsity on ViT-B/32.
-(a) Training coefficients only
-(b) Joint training
-Figure H: Impact of joint training on coefficient sparsity on ViT-B/16.
-(a) Training coefficients only
-(b) Joint training
-Figure I: Impact of joint training on coefficient sparsity on ViT-L/14.
 D.4 Sparsity Visualization
-The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging as shown in Figure 3.
+The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging.
 To confirm this trend across various backbones, Figure GH and I provide the layer-wise merging coefficients for ViT-B/32, ViT-B/16 and ViT-L/14.
@@
 These results indicate that training the classifier complements sparsity in merging coefficients, effectively reducing task conflicts and enhancing performance.
+(a) Training coefficients only
+(b) Joint training
+Figure G: Impact of joint training on merging coefficient sparsity on ViT-B/32.
+(a) Training coefficients only
+(b) Joint training
+Figure H: Impact of joint training on merging coefficient sparsity on ViT-B/16.
+(a) Training coefficients only
+(b) Joint training
+Figure I: Impact of joint training on merging coefficient sparsity on ViT-L/14.